### v3

# Combining Constrained Diffusion Models and Numerical Solvers for Efficient and Robust Non-Convex Trajectory Optimization

Abstract
Motivated by the need to solve open-loop optimal control problems with computational efficiency and reliable constraint satisfaction, we introduce a general framework that combines diffusion models and numerical optimization solvers. Optimal control problems are rarely solvable in closed form, hence they are often transcribed into numerical trajectory optimization problems, which then require initial guesses. These initial guesses are supplied in our framework by diffusion models. To mitigate the effect of samples that violate the problem constraints, we develop a novel constrained diffusion model to approximate the true distribution of locally optimal solutions with an additional constraint violation loss in training. To further enhance the robustness, the diffusion samples as initial guesses are fed to the numerical solver to refine and derive final optimal (and hence feasible) solutions. Experimental evaluations on three tasks verify the improved constraint satisfaction and computational efficiency with 4$\times$ to 30$\times$ acceleration using our proposed framework, which generalizes across trajectory optimization problems and scales well with problem complexity.
1 Introduction
Optimal control problems are fundamental in decision-making for autonomous agents including self-driving cars, robot manipulators, and spacecraft trajectories to name just a few. In the open-loop case, with a predetermined initial configuration, the goal is to plan an optimal path and control history for an agent to reach a target while satisfying the dynamics and safety constraints. These problems often have non-convex structure, which may be due to complex nonlinear system dynamics and environmental constraints. Ultimately, this results in multiple minima and solutions that may have different qualitative behavior with various tradeoffs. Identifying these diverse and locally optimal (and hence feasible) solutions efficiently remains a significant challenge in the field.
The open-loop optimal control problem can be cast as a numerical trajectory optimization problem using a variety of approaches [4]. When the optimal control problem involves nonlinear functions, this often results in a nonlinear program (NLP) where an initial guess is required to get the solution. In the case where one is interested in solving for several solutions, a global search is needed. This is essentially a two-step process: (i). sampling a distribution on the control space, and (ii). using this sample as an initial guess to the NLP solver that hopefully generates a solution. For each solution from the NLP, optimality (and hence feasibility) is often verified by the solver by checking whether the first order necessary conditions, the Karush-Kuhn-Tucker (KKT) conditions, are satisfied [29, 25]. However, the high dimensionality and non-convex nature of the problem often make the two-step global search process time-consuming and cannot satisfy the efficiency demands of autonomous systems.
Figure 1: Three trajectory optimization tasks in our experiments (from left to right): tabletop manipulation (161 DoF); two-car reach-avoid problem (81 DoF); cislunar transfer problem (64 DoF). The degree-of-freedom (DoF) indicates the dimension of control variables for each task.
Machine learning has also gained popularity in trajectory optimization, like warm starting the convexified trajectory optimization problems such as Quadratic Programming (QP) [11, 9, 38]. These approaches primarily focus on predicting a single solution to the QP problem to improve the computational efficiency, rather than modeling the distribution of diverse solutions to the original non-convex problem. Recently, generative models, particularly diffusion models, have shown good performance in generating trajectories or controls from a distribution [24, 1, 12, 8, 42]. When conditioned on the environmental parameters, diffusion models can efficiently generate appropriate and diverse trajectories even in a priori unknown environment. However, the diffusion model output trajectories are not robust to prediction error and may violate the dynamic and safety constraints, which can cause catastrophic damage to safety-critical systems. It’s also difficult for the diffusion model to directly output locally optimal trajectories. While efforts have been made to integrate safety constraints into generative models [32, 10, 49, 48, 36, 6], it is still challenging to achieve solution quality, computational efficiency, diversity, robustness and feasibility.
In this paper, we propose a general framework for efficient non-convex trajectory optimization with enhanced robustness for constraint satisfaction. This framework creates a synergy between a novel constrained diffusion model and a numerical optimization solver. We first formulate the trajectory optimization problem as a nonlinear program, where a numerical solver can be used to collect solutions as training data. Then we train a conditional diffusion model to efficiently sample a set of diverse and high-quality trajectory predictions, from which the solver can fine-tune to obtain final solutions with enhanced feasibility and optimality. We also developed an innovative constrained diffusion model that incorporates an additional constraint violation loss in the training process to minimize the violation in the samples.
Our framework has the following appealing properties: (a). Computational Efficiency. Trained with offline solution data, the diffusion model can predict high-quality solutions that are close to local optima and warm-start the trajectory optimization. This significantly accelerates the solver convergence with considerable computational time savings. (b). Robustness (Constraint Satisfaction). We used a constrained diffusion model specifically designed to produce samples with minimal constraint violations. These sampled trajectories are then further refined by a numerical solver to enhance the feasibility and optimality. (c). Automatic Data Generation. With a numerical solver, we can automatically generate diverse problems and collect corresponding solutions, which allows for on-demand data augmentation whenever required. (d). Generalization. Our framework is suitable for trajectory optimization across various applications with flexible environment settings.
We demonstrate the efficacy of our framework on three non-convex trajectory optimization problems for open-loop optimal control: tabletop manipulation, two-car reach-avoid, and cislunar low thrust transfer for a spacecraft. Our framework improves the efficiency of obtaining solutions compared to the baseline while enhancing constraint satisfaction.
2 Related Work
Machine Learning for Optimization.
Machine learning methods have been widely used to solve a variety of optimization problems with different techniques, including convex Quadratic Programming (QP) [50, 51, 11, 38], mixed-integer programming [9], combinatorial optimization [43, 23] and black-box optimization [30, 45, 28], etc. To improve the constraint satisfaction in optimization problems, neural networks have been incorporated with fully differentiable constraint function [16]. Conditional variational autoencoder (CVAE) and long short-term memory (LSTM) are proposed to warm-start the general nonlinear trajectory optimization problems [32]. Another branch of work leverages machine learning methods and pre-collected solution data points for solving partial differential equations (PDEs) [7]. One of the most pupular work is physics-informed neural networks (PINN) [37], which introduces the solution prediction model with a hybrid training loss for labeled solution prediction and violation of the original PDEs.
Diffusion Models.
The diffusion models [39, 41] is able to sample the multi-modal distribution with a stochastic differential equation (SDE). The models are further developed into denoising diffusion probabilistic models (DDPM) [21], denoising diffusion implicit models (DDIM) [40] for application of image generation. Both classifier-guided diffusion [13] and classifier-free guidance [22] are proposed to achieve conditional generation with diffusion models. Given the strong expressiveness of the diffusion model, its application has been extended across domains from image generation to reinforcement learning [46, 14, 15], motion generation [44], robotics [12], etc.
Diffusion-based Trajectory Generation.
Diffusion models are becoming popular in robotics for generating controls or trajectories [1, 24, 33, 42, 12, 34, 15]. Many efforts have been made to improve the diffusion model for creating safe trajectories for safety-critical systems. Control barrier functions [48, 6] and collision-avoidance kernels [10] are employed to guide diffusion models to sample feasible trajectory solutions. To adapt to new constraints in testing, prior work investigated composing diffusion models to improve the generalizability [36, 49]. However, solely relying on diffusion-sampled trajectories will inevitably face the robustness issue of constraint violations and hardly achieve locally optimal solutions. There has been an effort to combine diffusion models with an inference-based trajectory optimization framework [36, 35], but it is not parallelizable and requires extensive computational effort for complex problems. Our work guides the diffusion model toward low constraint violation in the output through a novel loss function and combines it with a fully parallelizable numerical optimization method that scales well.
3 Preliminaries
Non-Convex Trajectory Optimization.
We consider a general open-loop optimal control problem: assume a known agent’s dynamical model, the predefined initial configuration, and the obstacle settings, we aim to plan an optimal trajectory and controls for the agent to reach a target while satisfying the dynamics and safety constraints. With the direct transcription [4], this trajectory optimization problem can be discretized as the following nonlinear program $\mathcal{P}_{y}$:
[…]
where $x\in\mathcal{X}\subset\mathbb{R}^{n}$ is the decision variable to be optimized, and $y\in\mathcal{Y}\subset\mathbb{R}^{k}$ represents various problem parameters. $J\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ is the objective function. Each $h_{i}\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ and $g_{j}\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ are equality and inequality constraint functions, respectively. In trajectory optimization, $x$ might include a sequence of control $u$, the time variable $t$ required to reach the target, etc. Parameters $y$ can include control limits, target and obstacle characteristics such as position and shape, and other relevant factors. The objective function $J$ can be time-to-reach, fuel expenditure, etc. The constraint functions $g$ and $h$ could contain requirements for dynamical models, goal-reaching, obstacle avoidance, etc.
When the problem parameter $y$ is fixed, $\mathcal{P}_{y}$ in Eq. (1) typically presents a non-convex optimization problem with multiple local optima. To solve this problem $\mathcal{P}_{y}$, a gradient-based numerical solver $\pi$ can be used. A local optimum $x^{*}$ to $\mathcal{P}_{y}$ is verified by the solver $\pi$ that satisfies the KKT conditions (i.e., no constraint violation). To find different $x^{*}$, the user first provides multiple initial guesses $x^{0}$ for the decision variables $x$ to the solver $\pi$, typically drawn from a uniform distribution or some heuristic approach. Then the solver $\pi$ performs an iterative, gradient-based optimization until a local optimum $x^{*}$ is found. The popular methods to conduct such gradient-based optimization in non-convex problems include the Interior Point Method (IPM) [47] and the Sequential Quadratic Programming (SQP) [5].
Diffusion Probabilistic Model.
As expressive generative models, the diffusion probabilistic models [39, 21, 41] generate samples from random noises through an iterative denoising process. The forward process (i.e., diffusion process) iteratively adds Gaussian noise to sample $x_{k}$ at the $k$-th step according to the following equation:
$q(x_{k+1}|x_{k})=\mathcal{N}(x_{k+1};\sqrt{1-\beta_{k}}x_{k},\beta_{k}\mathbf{I}),\quad 0\leq k\leq K-1.$ (2)
where $\beta_{k}$ is a given variance schedule. With the diffusion steps $K\rightarrow\infty$, $x_{K}$ becomes a random Gaussian noise. From above one can write:
$x_{k}=\sqrt{\widebar{\alpha}_{k}}x_{0}+\sqrt{1-\widebar{\alpha}_{k}}\varepsilon,$ (3)
where $\widebar{\alpha}_{k}=\prod_{k^{\prime}=1}^{K}(1-\beta_{k^{\prime}})$ and $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ is a random sample from a standard Gaussian. The data generation process is the reverse denoising process that transforms random noise to data sample:
$p_{\theta}(x_{k-1}|x_{k})=\mathcal{N}\big{(}x_{k-1};\mu_{\theta}(x_{k}),\sigma_{k}\mathbf{I}\big{)},\quad 1\leq k\leq K$ (4)
with initial noisy data sampled from a random Gaussian distribution $x_{K}\sim\mathcal{N}(0,\mathbf{I})$. $p_{\theta}$ is the parameterized sampling distribution, and it can be optimized through an alternative $\varepsilon_{\theta}(x_{k},k)$ which is to predict $\varepsilon$ injected for $x_{k}$, for every diffusion step $k$. For conditional generation modeling, a conditional variable $y$ is added to both processes $q(x_{k+1}|x_{k},y)$ and $p_{\theta}(x_{k-1}|x_{k},y)$, respectively. The classifier-free guidance [22] can be applied to further promote the conditional information, which learns both conditioned and unconditioned noise predictors as $\varepsilon_{\theta}(x_{k},k,y)$ and $\varepsilon_{\theta}(x_{k},k,\varnothing)$. Specifically, $\varepsilon_{\theta}$ is optimized with the following loss function:
$\mathcal{L}_{\text{diff}}=\operatorname{\mathbb{E}}_{(x_{0},y)\sim\mathcal{X}\times\mathcal{Y},k,\varepsilon,b}\left\|\varepsilon_{\theta}\Big{(}x_{k}(x_{0},\varepsilon),k,(1-b)\cdot y+b\cdot\varnothing\Big{)}-\varepsilon\right\|^{2}_{2}$ (5)
where $x_{0}$ and $y$ are sampled from groundtruth data, $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$, $k\sim\text{Uniform}(1,K)$, $b\sim\text{Bernoulli}(p_{\text{uncond}})$ with given unconditioned probability $p_{\text{uncond}}$, and $x_{k}$ follows Eq. (3).
4 Methodology
In this section, we first proposed the general framework for efficient and robust non-convex trajectory optimization using diffusion models and numerical solvers. Then we introduce a novel constrained diffusion model specifically designed to guide the diffusion models to generate feasible solutions.
4.1 A General Trajectory Optimization Framework
Figure 2: The proposed general framework for non-convex trajectory optimization with a constrained diffusion model and a numerical solver.
It has been observed that for similar trajectory optimization problems $\mathcal{P}_{y}$ in Eq. (1), their solutions $x^{*}$ often exhibit similar structures [2, 32]. Similar problems are those $\mathcal{P}_{y}$ with the same objective $J$ and constraints $g,h$, but with different parameters $y$. Thus, the key insight of our framework is to use a diffusion model to learn $p(x^{*}|y)$, the conditional probability distribution of locally optimal solutions, based on pre-solved similar problems. Then the diffusion model can generalize and predict diverse solutions $\tilde{x}_{\text{new}}^{*}$ to new scenarios where the condition $y_{\text{new}}$ value is unseen.
However, diffusion models inevitably make prediction errors, which can result in constraint violations. To address this, we use diffusion samples $\tilde{x}_{\text{new}}^{*}$ as initial guesses for the numerical solver, which fine-tunes them to derive the final solutions $x_{\text{new}}^{*}$, increasing both feasibility and optimality. This approach acts like a safety net, enhancing the robustness of the solutions.
We illustrate our proposed data-driven trajectory optimization framework with diffusion models in Fig. 2. In the offline data generation process, we use a numerical solver $\pi$ to collect locally optimal solutions $x^{*}$ for various problem instances $\mathcal{P}_{y}$ with uniformly sampled initial guesses. The solutions with sub-optimal objective values are filtered out manually and the remaining dataset is used as the training data. We train a conditional diffusion model to learn $p(x^{*}|y)$ with classifier-free guidance [22]. In the online testing process when a new problem $\mathcal{P}_{y_{\text{new}}}$ is presented with a priori unknown value $y_{\text{new}}$, we first use the diffusion model to predict diverse solution candidates $\tilde{x}_{\text{new}}^{*}$ as initial guesses. Then the numerical solver $\pi$ only requires minor adjustments to derive the final solutions $x_{\text{new}}^{*}$ with improved feasibility and optimality. This warm-starting approach is fully parallelizable for both diffusion sampling and the solving process of $\pi$.
4.2 Constrained Diffusion Model
The original diffusion models (DMs) aim to learn the distribution of locally optimal solutions $p(x^{*}|y)$ through the loss function in Eq. (5) without knowing any constraint information. As a result, the sampled solutions could be close to the groundtruth but still largely violate the constraints. To address this issue, we propose a novel constrained diffusion model (CDM) to minimize the constraint violations in the sampled trajectories. Inspired by the equation loss in PINN [37], we define a violation function $V(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for differentiable constraints in Eq. (1):
$V=\sum_{i=1}^{l}\max(g_{i},0)+\sum_{j=1}^{m}|h_{j}|,$ (6)
where $V(\cdot,\cdot)\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ (differentiable except at $g_{i}=0$ or $h_{j}=0$) maps from a sample $x$ and parameter $y$ and outputs the total constraint violation as a scalar value. In fact, each constraint term can be further customized with different weights that allow different treatments. The violation value is expected to decrease during the training process of diffusion models. We introduce the newly proposed loss function and the corresponding training process as follows.
CDM Training.
We define the locally optimal solution $x^{*}=x_{0}$ as initial sample in the diffusion process with diffusion step $k=0$. By reformulating the forward sampling process as in Eq. (3), we have $x_{0}=\frac{x_{k}-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon}{\sqrt{\widebar{\alpha}_{k}}},\varepsilon\sim\mathcal{N}(0,\mathbf{I})$. The noise $\varepsilon$ is approximated as $\varepsilon_{\theta}(x_{k},k,y)$ with neural-network parameterization. Therefore, we can predict initial sample as $\widehat{x}_{0}=\frac{x_{k}(x_{0},\varepsilon)-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon_{\theta}(x_{k},k,y)}{\sqrt{\widebar{\alpha}_{k}}}$. Then we introduce a constraint violation loss $\mathcal{L}_{\text{vio}}$ on the predicted $\widehat{x}_{0}$ from $x_{k}$, $\forall k\in[K]$:
$\mathcal{L}_{\text{vio}}$ $=\mathbb{E}_{(x_{0},y)\sim\mathcal{X}\times\mathcal{Y},k\in[K],\varepsilon\sim\mathcal{N}(0,\mathbf{I})}\bigg{[}\frac{1}{k}V\Big{(}\text{clip}\big{(}\widehat{x}_{0}(x_{k},\varepsilon_{\theta}),x_{0}^{\text{min}},x_{0}^{\text{max}}\big{)},y\Big{)}\bigg{]}$ (7)
where $x_{0}^{\text{min}}$ and $x_{0}^{\text{max}}$ are the lower and upper bound for $x_{0}\in\mathcal{X}$. The reason for the clipping is that at the early phase of the training, the parameterized $\varepsilon_{\theta}$ as an approximation of $\varepsilon$ is not accurate thus the predicted $\widehat{x}_{0}$ may not lie on the original range of the data $x_{0}$. Therefore, we clip $\widehat{x}_{0}$ to maintain its range, which guarantees the valid calculation of violation function $V(\cdot,\cdot)$. The scheduling factor $\frac{1}{k}$ regulates the scale of the violation function $V$ for different $x_{k},k\in[K]$. The reason is that the predicted $\widehat{x}_{0}$ will be more noisy when $k$ is large therefore the estimation of violation function can be less reliable. This may not be a theoretically principled choice but is found effective and convenient in our experiments.
For the violation loss $\mathcal{L}_{\text{vio}}$, we always let the condition variable $y$ appear in the noise prediction $\varepsilon_{\theta}(x_{k},k,y)$ without masking it out, contrary to the probabilistic masking in the original training loss $\mathcal{L}_{\text{diff}}$. An intuitive interpretation of the additional loss term is that it guides the diffusion model training with the gradients of minimizing the violation value for each denoising step.
Finally, our proposed constrained diffusion training loss is a combination of the original diffusion model training loss and the above violation loss:
$\mathcal{L}=\mathcal{L}_{\text{diff}}+\lambda\mathcal{L}_{\text{vio}}$ (8)
where $\lambda$ is a hyperparameter for adjusting the strength of $\mathcal{L}_{\text{vio}}$. As the training proceeds, the original loss function $\mathcal{L}_{\text{diff}}$ will encourage $\widehat{x}_{0}$ to be close to true $x_{0}$, and $\mathcal{L}_{\text{vio}}$ further enforces the constraint satisfaction of the generated samples.
Conditional Diffusion Sampling.
For sampling from trained diffusion models, we adopt guided sampling with classifier-free guidance as introduced in Sec. 3. The guided prediction noise with guidance weight $c$ is:
$\widehat{\varepsilon}_{\theta}\leftarrow c\cdot\varepsilon_{\theta}(x_{k},k,y)+(1-c)\cdot\varepsilon_{\theta}(x_{k},k,\varnothing)$ (9)
Then we plug $\widehat{\varepsilon}_{\theta}$ into $\varepsilon_{\theta}$ in the sampling process [21, 22], and through iterative sampling $x_{k-1}$ from $x_{k}(k=K,\dots,1)$ with initial random noise $x_{K}\sim\mathcal{N}(0,\mathbf{I})$, it produces the generated samples $x_{0}$. This diffusion samples serve as the predicted initial guess in Sec. 4.1 for a new task: $\tilde{x}^{*}_{\text{new}}:=x_{0}$.
5 Experiment
In this section, we evaluate the proposed framework combining diffusion models and numerical solvers on three trajectory optimization tasks, and compare with several baselines. Our methods include two DMs: the vanilla diffusion model (DM) and the constrained diffusion model (CDM) as proposed in Sec. 4.2. The experiments involve three stages: 1. Collect solutions provided by numerical solvers on each task domain; 2. For data-driven methods, train the models to fit the solution dataset; 3. For unseen new tasks within each domain, sample initial solutions from models and feed into numerical solvers for deriving final solutions. The details of data collection and model training refer to Appendix A.
Tasks.
Our experiments are conducted on three trajectory optimization tasks across different domains with examples, as shown in Fig. 1. With an increasing complexity (non-convexity) of the problem, we introduce: (i). tabletop manipulation for object relocation with a robotic arm (Sec. 5.1); (ii). the two-car reach-avoid problem as a two-player motion planning game (Sec. 5.2); and (iii) the cislunar transfer problem for optimal control of the spacecraft (Sec. 5.3). Each task has various DoF, which is derived with the number of control variables at each time step multiplied with the number of discretized steps along the trajectory. The tabletop manipulation has a higher DoF due to a larger number of time stesp ($T=80$), however, the task is less complex as the cislunar transfer, which has more complicated dynamics thus higher non-convexity. More details for each task refer to Appendix A.
Numerical Solvers.
We assume fully known dynamics and environment for the open-loop optimal control problem. To solve the tabletop manipulation and two-car reach-avoid problem, we formulate and solve the problem with Sparse Nonlinear OPTimizer (SNOPT) [19] as the solver $\pi$, which uses the SQP method and supports warm-starting. For the cislunar transfer, we formulate and solve the problem with pydylan, a Python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] together with SNOPT. The detailed problem formulation and solving for each task are included in Appendix A.
Baselines.
For comparison with the proposed framework, we introduce the following baselines:
  • Uniform: This method solely uses a numerical solver with uniformly sampled initial guesses $x\in\mathcal{X}$ for the non-convex trajectory optimization problem. It solves problems from scratch without leveraging any problem-specific prior information.
  • Optimal Uniform: This method uniformly samples from collected locally optimal solutions $\{x^{*}\}$ as initial guesses for the numerical solver, without using any machine learning method to approximate the solution distribution.
  • CVAE_LSTM: This method [32] combines the convolutional variational auto-encoder (CVAE) with long short-term memory (LSTM) to learn the conditional distribution of locally optimal solutions. Specifically, a CVAE with Gaussian Mixture Model (GMM) prior is first used to sample the non-control variables, e.g. time or mass variables, and then an LSTM is adapted to sample the control sequence.
Although there exist other graph-based path planning methods like Rapidly-exploring Random Trees (RRT) [31], A [20], etc., they can not account for both complex dynamical and environment constraints and flexible optimization objectives in the optimal control problem. Consequently, the comparison against them is not provided in our experiments.
Diversity Measure.
Apart from the time consumption for deriving feasible solutions, we also apply diversity measure for each method since there exist multiple local optima for each task. Specifically, the Vendi Score [18] is applied to measure the diversity of the model outputs. The similarity kernel we choose is $k(x,y)=e^{-||x-y||}$.
5.1 Tabletop Manipulation
Task Setting.
For a robotic arm manipulating objects on the tabletop as shown in Fig. 7 in the Appendix, it is important to plan a desired trajectory with goal-reaching and obstacle avoidance. We define the states of the system to be the location of the gripper: $(p_{x},p_{y})\in[0,10]^{2}$. In a simplified simulation setting, the movement of the gripper is modeled with linear dynamics $\dot{p}_{x}=v_{x},\dot{p}_{y}=v_{y}$, where $u=(v_{x},v_{y}),v_{x},v_{y}\in[-1,1]$ are two-dimensional speed control variable. In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles.
Figure 3: Raw trajectory prediction output and the solver output for the tabletop manipulation (top row) and two-car reach-avoid problem (bottom row). From left to right, uniform sampling, CVAE_LSTM, DM, CDM and solver refined locally optimal solution from CDM sampled initial guess.
Results.
In the top row of Fig. 3, our proposed constrained diffuse model generates a nearly collision-free trajectory to reach the goal, which is very similar to the final refined solution from the solver. The DM tries to navigate through the cluttered environment but chooses a path that inevitably collides with obstacles. The baseline CVAE_LSTM is not able to account for both goal-reaching and obstacle avoidance. This can be further verified by checking the constraint violation in Table 2, where both DM and CDM achieve significant reduction on constraint violation. CDM has the lowest violation, which justifies the effectiveness of the additional constraint violation loss for training.
When we use our proposed models to warm-start the solver $\pi$, only minor refinement is needed to derive locally optimal trajectories. We compute the ratio of locally optimal solutions obtained from 600 initial guesses and present the corresponding time consumption (involving both sampling time from the model and numerical solver time) statistics in Table 1. As demonstrated in Table 1, the CDM with the solver is 29.6 times faster than the uniform method based on the top $25\%$-quantile of the computational time for deriving the locally optimal solutions. This is further demonstrated in Fig. LABEL:Fig:tabletop_time_histogram with the solving time histogram of locally optimal solutions, where the CDM with the solver starts to obtain many solutions in a shorter solving time. Then Vendi score result in Table 1 verifies the diversity of our model output, where uniform samples intrinsically have the highest diversity. Thus both efficiency and solution quality are achieved with our framework.
5.2 Two-Car Reach-Avoid
Task settings.
The two-car reach-avoid problem is a two-player game where each car tries to find a minimum-time path to reach the goal while avoiding collision with another car and obstacles simultaneously. Specifically, both car1 and car2 follows the same dynamics as follows:
$\dot{p}_{x}=v\cos{\theta},\quad\dot{p}_{y}=v\sin{\theta},\quad\dot{v}=a,\quad\dot{\theta}=\omega,\quad$
where $(p_{x},p_{y},v,\theta)$ denotes the state of each car. $p_{x},p_{y}$ are the position, $v$ is the speed, and $\theta$ is the orientation of the car. Each car has two-dimensional controls $u^{1}=(a^{1},\omega^{1}),u^{2}=(a^{2},\omega^{2})$, where $a\in[-1,1]$ is the acceleration and $\omega\in[-1,1]$ is the angular speed. The task setup can be seen in Fig. 1.
Table 1: The ratio of locally optimal solutions, the statistics (mean, standard deviation, quantile and median) of total time consumption including both model sampling time and solver time, and Vendi scores for diversity of sampled trajectories before feeding to the solver.
Task Method Ratio Mean($\pm$±STD) 25%-Quantile Median Vendi Score
Tabletop CDM (Ours) 59% 7.29 $\pm$ 16.57 0.50 1.26 5190.70
DM (Ours) 60% 7.51 $\pm$ 17.50 0.73 1.37 5306.18
CVAE_LSTM 53% 8.53 $\pm$ 20.87 1.10 1.55 516.61
Uniform 62% 31.65 $\pm$ 24.55 14.81 22.72 5917.73
Optimal Uniform 77% 12.80 $\pm$ 17.70 4.11 7.17 3708.58
Two-car CDM (Ours) 95% 17.86 $\pm$ 14.28 7.68 13.83 3873.60
DM (Ours) 95% 18.82 $\pm$ 14.33 8.77 15.61 3574.66
CVAE_LSTM 63% 36.24 $\pm$ 20.55 19.45 33.26 1162.72
Uniform 64% 46.17 $\pm$ 20.63 29.62 43.54 5256.75
Optimal Uniform 83% 25.15 $\pm$ 19.75 10.82 20.89 3708.58
Cislunar DM (Ours) 54% 50.81 $\pm$ 70.30 9.97 26.68 5999.96
CVAE_LSTM 30% 141.31 $\pm$ 125.60 38.41 106.48 1403.38
Uniform 18% 199.34 $\pm$ 116.59 114.02 174.35 6000
Optimal Uniform 23% 177.35 $\pm$ 123.22 70.03 153.97 5918.08
Table 2: The constraint violation value for generated samples before feeding to the solver.
Task Method Mean($\pm$±STD)
Tabletop CDM (Ours) 4.60 $\pm$ 5.38
DM (Ours) 4.73 $\pm$ 5.73
CVAE_LSTM 9.11 $\pm$ 8.16
Uniform 32.20 $\pm$ 3.73
Optimal Uniform 65.44 $\pm$ 40.22
Two-car CDM (Ours) 3.00 $\pm$ 15.36
DM (Ours) 10.99 $\pm$ 35.59
CVAE_LSTM 282.72 $\pm$ 186.40
Uniform 545.66 $\pm$ 294.71
Optimal Uniform 3.27 $\pm$ 3.57
Results.
In the bottom row of Fig. 3, the proposed CDM generates the best prediction of the trajectory solution conditioned on obstacle information $y$, which is very close to the final locally optimal solutions refined by the solver. The unconstrained DM samples reasonable trajectory solution prediction, but with much higher constraint violation. The baseline CVAE_LSTM cannot generate solutions toward goal-reaching. Uniformly sampled initial guesses are usually far from a constraint-satisfying trajectory.
We also evaluate the solving time of obtaining those locally optimal solutions $x^{*}$ from the solver $\pi$ for these five methods in Table 1, and visualize the corresponding histogram in Fig. LABEL:Fig:car_solving_time_histogram. Our proposed CDMs obtained the greatest number of locally optimal solutions from 600 samples as initializations. It is also 3.9 times faster than the uniform method according to the $25\%$-quantile of the computational time for deriving the locally optimal solutions.
To further compare the CDM and the unconstrained DM, we also measure the constraint violation (as defined by Eq. (6)) of all the model outputs in Table 2. Our CDM achieves the lowest constraint violation among all data-driven models. Note that the optimal uniform also achieves small violation because in this task the goal conditions are fixed, the pre-collected solutions will only incur obstacle collision costs which are distance to the obstacle boundary and are inherently small.
The Vendi score in Table 1 shows the superiority of using diffusion models to generate diverse outputs compared to the baseline CVAE_LSTM. We also visualize the qualitatively different trajectory solution with the proposed CDM and the corresponding solver in Figure 4.
Figure 4: Visualization of four solutions for one two-car reach-avoid problem, with CDM predicted initial guesses refined by the solver $\pi$.
5.3 Cislunar Transfer
Task settings.
In this task, we consider a minimum-fuel cislunar transfer with the dynamical model to be a Circular Restricted Three-Body Problem (CR3BP) [27]. Assuming the mass of the spacecraft is negligible, CR3BP describes the equation of motions of a spacecraft under the gravitational force from the sun and moon. Let $m_{1}$ be the mass of the Earth and $m_{2}$ be the mass of the moon, and $\mu=m_{2}/(m_{1}+m_{2})$, the dynamics of the spacecraft in the CR3BP is as follows:
$\ddot{x}-2\dot{y}=-\widebar{U}_{x},\quad\ddot{y}+2\dot{x}=-\widebar{U}_{y},\quad\ddot{z}=-\widebar{U}_{z},$
where $\widebar{U}(r_{1}(x,y,z),r_{2}(x,y,z))=-\frac{1}{2}\left((1-\mu)r_{1}^{2}+\mu r_{2}^{2}\right)-\frac{1-\mu}{r_{1}}-\frac{\mu}{r_{2}}$ represents the effective gravitational potential, $r_{1}=\sqrt{(x+\mu)^{2}+y^{2}+z^{2})}$ and $r_{2}=\sqrt{(x-(1-\mu))^{2}+y^{2}+z^{2})}$ are the distance from the spacecraft to the sun and moon, respectively. The objective is to minimize fuel expenditure.
Results.
The cislunar transfer is a challenging problem to solve that requires intensive computational time. In Table 1, our method with the unconstrained DM has a huge improvement in the average solving time and the number of locally optimal solutions obtained, compared to the baseline. It is also 7.0 times faster than the uniform method based on the $25\%$-quantile of computational time for deriving the locally optimal solutions. This can be further verified by the solving time histogram in Fig. LABEL:Fig:_cr3bp_solving_time_histogram. This improvement encourages us to set a short cut-off time for the solver and use more time to sample different initial guesses, which further improves the efficiency of obtaining solutions. Finally, the Vendi score shown in Table 1 indicates that our unconstrained DM generates much more diverse predictions than CVAE_LSTM. Since the cislunar transfer task uses a complex numerical integration method for the dynamics constraint which requires extra effort to incorporate it into the neural network, we will leave the CDM for this task as future work.
6 Conclusion and Discussion
This paper presents a general and parallelizable framework for non-convex trajectory optimization, using a novel constrained diffusion model combined with numerical solvers. It efficiently provides divesrse solutions for problems with mutiple local optima, with less constraint violation, and up to 30 times acceleration. It is a general acceleration framework for solving scientific optimization problems.
Limitations also exist for the current framework. Although most constraint functions in trajectory optimization are differentiable (almost everywhere), current CDM does not work for non-differentiable violation functions. For non-differentiable constraints, transforming them into differentiable functions or applying approximation with appropriate bump functions or neural networks is possible. Another problem is that the diffusion sampling may not be achieved in real time in a highly dynamic environment, which can be improved with stride sampling, DDIM [40], etc.
Appendix A Implementation and Task Details
A.1 Model Details
For both unconstrained and constrained diffusion models, we apply a UNet structure with three hidden layers of 512, 512 and 1024 neurons. We also use a fully connected layer of 256 and 512 neurons to embed the conditional input $y$. The sampling step is set to be 500. Both DMs are trained with 200 epochs using the Adam optimizer [26]. The constrained and unconstrained DMs usually take 14 hours and 6 hours respectively, as well as 20 GB VRAM to train on one NVIDIA A100 GPU, respectively.
A.2 Tabletop Manipulation
Problem formulation and solving.
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles. The start position is fixed at $(5,5)$. The conditional parameter $y$ includes the goal position $p_{\text{goal}}$ randomly chosen from $\{(1,1),(1,9),(9,1),(9,9)\}$, 4 obstacle positions $p_{\text{obs}}$ randomly sampled between the start and goal, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.0]$. We discretize this problem into $T$ timesteps where $T=80$. The variable to optimize is $x=(t_{\text{final}},u_{1},u_{2},..,u_{T})\in\mathbb{R}^{161}$, and the conditional parameter is $y=(p_{\text{goal}},p_{\text{obs}},r_{\text{obs}})\in\mathbb{R}^{14}$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$ and the collision avoidance for obstacles at all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
Data collection and model training.
In this task, to collect the training data we sample 2700 different $y$, each with 100 uniformly sampled initial guesses. Then we use the SNOPT [19] solver for each problem instance to collect locally optimal solutions. Finally, we filter out those solutions with objective values greater than 6 s and use the remaining 237k data as training data. Within this 237k data, 10% of the data are used as the validation set. The whole data collection process takes around 15 hours with 200 AMD EPYC 9654 CPU cores.
We train each of our proposed constrained and unconstrained DMs and baseline models with 3 different random seeds. The models are tested for the warm-starting performance with 600 initial guesses from 3 random seeds across 20 unseen locations of the goal and obstacles, i.e. unseen $y$ values.
[…]
Figure 7: An example of tabletop manipulation task for goal-reaching and obstacle avoidance movement, using a CDM with a numerical solver for planning: (left) starting position of the gripper and the planned trajectory; (right) target position reached by the gripper.
A.3 Two Car Reach-Avoid
Problem formulation and solving.
In this task $\mathcal{P}_{y}$, we aim to minimize the time for the two cars to reach each own goal while avoiding colliding with the other car and obstacles. We fixed the start $p_{\text{start}}$ and goal $p_{\text{goal}}$ position for car1 to be $(0,10)$ and $(10,0)$, and for car2 to be $(10,0)$ and $(0,0)$. The conditional parameters $y\in\mathbb{R}^{6}$ include the two obstacle positions $p_{\text{obs}}$ randomly sampled from $[2.0,8.0]$, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.5]$. We formulate this trajectory optimization problem with a forward shooting method, with the discretized time step $T=40$. The variable to optimize is $x=(t_{\text{final}},u^{1}_{1},u^{1}_{2},...,u^{1}_{T},u^{2}_{1},u^{2}_{2},...,u^{2}_{T})$, where $u^{1}_{i}$ and $u^{2}_{i}$ are controls for two cars, respectively. The problem parameter is $y=(p_{\text{obs}},r_{\text{obs}})$. The objective is defined as $J(x;y)=t_{\text{final}}$. The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$, and the collision avoidance for obstacles and between each car for all time. We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
Data collection and model training.
In this task, we collect locally optimal solutions $x^{*}$ from 3k different $y$, each with 100 uniformly sampled initial guesses using SNOPT [19]. We filter out those solutions whose objective is higher than 12 s and use the remaining 114k solutions as the training data. Within this 114k data, 10% of the data are used as the validation set. The whole data collection process takes around 8 hours with 200 AMD EPYC 9654 CPU cores.
We train each of our proposed constrained and unconstrained DMs and baseline models with 3 different random seeds and test the warm-starting performance with 600 initial guesses sampled from 3 random seeds across 20 unseen obstacle settings, i.e. unseen $y$ values.
A.4 Cislunar Transfer
Problem formulation and solving.
In this problem $\mathcal{P}_{y}$, as shown in Fig. 1, the spacecraft is planned to start from a Geostationary Transfer Spiral (blue) and reach a stable manifold arc (green) of a Halo orbit near $L_{1}$ Lagrange point (green). A candidate trajectory solution is plotted in orange. The CR3BP is a chaotic system, where a smaller perturbation will lead to a significantly different final trajectory. Therefore, trajectory optimization of this problem contains many local optima. We choose the conditional parameter $y\in[0.1,1.0]$ Newton to be the maximum allowable thrust of the spacecraft.
We use the forward-backward shooting method to formulate the trajectory optimization problem in Eq. (1). For the spacecraft, we set the constant specific impulse (CSI) $I_{sp}=1000$ s. The initial dry mass is 300 kg and the initial fuel mass is 700 kg. We choose $T$ discretized control segments where $T=20$ and the control $u$ are uniformly applied on each segment. The variable $x$ to optimize is as follows:
$\bm{x}=(t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}},m_{f},u_{1},u_{2},...,u_{N})$ (10)
where $t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}}$ are the effective burning time for the engine, initial and final coast time, $m_{f}$ is the final mass, and $u_{i}$ is a three-dimensional thrust control of the spacecraft. The objective is defined as $J(x;y)=-m_{f}$, which is to minimize fuel expenditure. The constraint $g$ is the midpoint agreement when integrating dynamics from forward and backward. The problem parameter $y$ is the maximum allowable thrust for the spacecraft and is bounded by $[0.1,1]$ Newton. To integrate the dynamics, we use the Runge-Kutta-Fehlberg 5-4th order (RK54) method [17].
Data collection and model training.
To collect the training data, we sample 12 different $y$ values from a grid in $[0.1,1]$ Newton, and for each $y$ we collect 25k locally optimal solutions with $m_{f}\geq 415kg$ total amount of 300k data. The problem is solved using pydylan, a python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] and the solver SNOPT [19]. Within this 300k data, 10% of the data are used as the validation set. The whole data collection process takes around 3 days with 500 Intel Skylake CPU cores.
We train the unconstrained DM and baseline models on 3 random seeds, and we test the warm-starting performance with 600 initial guesses sampled from 3 random seeds on $y=0.15$ Newton, which is unseen in training data.

### v4

@@
-Combining Constrained Diffusion Models and Numerical Solvers for Efficient and Robust Non-Convex Trajectory Optimization
+DiffuSolve: Diffusion-based Solver for Non-convex Trajectory Optimization
 Abstract
-Motivated by the need to solve open-loop optimal control problems with computational efficiency and reliable constraint satisfaction, we introduce a general framework that combines diffusion models and numerical optimization solvers.
-Optimal control problems are rarely solvable in closed form, hence they are often transcribed into numerical trajectory optimization problems, which then require initial guesses.
-These initial guesses are supplied in our framework by diffusion models.
-To mitigate the effect of samples that violate the problem constraints, we develop a novel constrained diffusion model to approximate the true distribution of locally optimal solutions with an additional constraint violation loss in training.
-To further enhance the robustness, the diffusion samples as initial guesses are fed to the numerical solver to refine and derive final optimal (and hence feasible) solutions.
-Experimental evaluations on three tasks verify the improved constraint satisfaction and computational efficiency with 4$\times$ to 30$\times$ acceleration using our proposed framework, which generalizes across trajectory optimization problems and scales well with problem complexity.
+Optimal trajectory design is computationally expensive for nonlinear and high-dimensional dynamical systems.
+The challenge arises from the non-convex nature of the optimization problem with multiple local optima, which usually requires a global search.
+Traditional numerical solvers struggle to find diverse solutions efficiently without appropriate initial guesses.
+In this paper, we introduce DiffuSolve, a general diffusion model-based solver for non-convex trajectory optimization.
+An expressive diffusion model is trained on pre-collected locally optimal solutions and efficiently samples initial guesses, which then warm-starts numerical solvers to fine-tune the feasibility and optimality.
+We also present DiffuSolve+, a novel constrained diffusion model with an additional loss in training that further reduces the problem constraint violations of diffusion samples.
+Experimental evaluations on three tasks verify the improved robustness, diversity, and a 2$\times$ to 11$\times$ increase in computational efficiency with our proposed method, which generalizes well to trajectory optimization problems of varying challenges.
 1 Introduction
-Optimal control problems are fundamental in decision-making for autonomous agents including self-driving cars, robot manipulators, and spacecraft trajectories to name just a few.
-In the open-loop case, with a predetermined initial configuration, the goal is to plan an optimal path and control history for an agent to reach a target while satisfying the dynamics and safety constraints.
-These problems often have non-convex structure, which may be due to complex nonlinear system dynamics and environmental constraints.
-Ultimately, this results in multiple minima and solutions that may have different qualitative behavior with various tradeoffs.
-Identifying these diverse and locally optimal (and hence feasible) solutions efficiently remains a significant challenge in the field.
-The open-loop optimal control problem can be cast as a numerical trajectory optimization problem using a variety of approaches [4].
-When the optimal control problem involves nonlinear functions, this often results in a nonlinear program (NLP) where an initial guess is required to get the solution.
-In the case where one is interested in solving for several solutions, a global search is needed.
-This is essentially a two-step process: (i).
-sampling a distribution on the control space, and (ii).
-using this sample as an initial guess to the NLP solver that hopefully generates a solution.
-For each solution from the NLP, optimality (and hence feasibility) is often verified by the solver by checking whether the first order necessary conditions, the Karush-Kuhn-Tucker (KKT) conditions, are satisfied [29, 25].
-However, the high dimensionality and non-convex nature of the problem often make the two-step global search process time-consuming and cannot satisfy the efficiency demands of autonomous systems.
-Figure 1: Three trajectory optimization tasks in our experiments (from left to right): tabletop manipulation (161 DoF); two-car reach-avoid problem (81 DoF); cislunar transfer problem (64 DoF).
-The degree-of-freedom (DoF) indicates the dimension of control variables for each task.
-Machine learning has also gained popularity in trajectory optimization, like warm starting the convexified trajectory optimization problems such as Quadratic Programming (QP) [11, 9, 38].
-These approaches primarily focus on predicting a single solution to the QP problem to improve the computational efficiency, rather than modeling the distribution of diverse solutions to the original non-convex problem.
-Recently, generative models, particularly diffusion models, have shown good performance in generating trajectories or controls from a distribution [24, 1, 12, 8, 42].
-When conditioned on the environmental parameters, diffusion models can efficiently generate appropriate and diverse trajectories even in a priori unknown environment.
-However, the diffusion model output trajectories are not robust to prediction error and may violate the dynamic and safety constraints, which can cause catastrophic damage to safety-critical systems.
-It’s also difficult for the diffusion model to directly output locally optimal trajectories.
-While efforts have been made to integrate safety constraints into generative models [32, 10, 49, 48, 36, 6], it is still challenging to achieve solution quality, computational efficiency, diversity, robustness and feasibility.
-In this paper, we propose a general framework for efficient non-convex trajectory optimization with enhanced robustness for constraint satisfaction.
-This framework creates a synergy between a novel constrained diffusion model and a numerical optimization solver.
-We first formulate the trajectory optimization problem as a nonlinear program, where a numerical solver can be used to collect solutions as training data.
-Then we train a conditional diffusion model to efficiently sample a set of diverse and high-quality trajectory predictions, from which the solver can fine-tune to obtain final solutions with enhanced feasibility and optimality.
-We also developed an innovative constrained diffusion model that incorporates an additional constraint violation loss in the training process to minimize the violation in the samples.
-Our framework has the following appealing properties: (a).
+Optimal trajectory design is fundamental in decision-making for autonomous agents.
+In the open-loop case, with a predetermined initial configuration, the goal is to plan an optimal path and control sequence for an agent to reach a target while satisfying the dynamics and safety constraints.
+In the case of complex nonlinear system dynamics and environmental constraints, these problems often have non-convex structures.
+For example, within a chaotic dynamical system like the Circular Restricted Three-Body Problem (CR3BP) (Koon et al., 2000), the trajectory optimization problem for the spacecraft has multiple local optima of different qualitative behaviors with various tradeoffs (Hartmann et al., 1998; Russell, 2007).
+Identifying these diverse, locally optimal, and feasible solutions efficiently remains a significant challenge in the field.
+Control transcription of the open-loop trajectory optimization problem involving nonlinear functions results in a nonlinear program (NLP), where an initial guess is required to get the solution (Betts, 1998).
+When one is interested in solving for several solutions, a two-step global search is needed: (i) sampling a distribution on the control space, and (ii) using this sample as an initial guess to the NLP solver that hopefully generates a solution.
+For each solution from the NLP, optimality (and also feasibility) is often verified by the solver by checking whether the first-order necessary conditions, the Karush-Kuhn-Tucker (KKT) conditions, are satisfied (Kuhn & Tucker, 2013; Karush, 1939).
+However, the high dimensionality and non-convex nature of the problem often make the two-step global search process time-consuming such that the efficiency demands of autonomous systems are not met.
+Figure 1: Three trajectory optimization tasks in our experiments (from left to right): two-car reach-avoid problem; quadrotor navigation problem; cislunar transfer problem.
+The dimensions of the decision variables are 81, 241, and 64, respectively.
+Machine learning has also gained popularity in trajectory optimization, but primarily focuses on solving the convexified problem rather than learning the solution distribution to the original non-convex problem.
+For example, investigations have been made on warm-starting Quadratic Programs (QP) using neural networks (Chen et al., 2022; Cauligi et al., 2021; Sambharya et al., 2023).
+Recently, generative models, particularly diffusion models, have shown good performance in generating trajectories or controls from a distribution (Janner et al., 2022; Ajay et al., 2022; Chi et al., 2023; Carvalho et al., 2023; Sridhar et al., 2023).
+When conditioned on the environmental parameters, diffusion models can efficiently generate appropriate and diverse trajectories even in an a priori unknown environment.
+However, the diffusion model alone without the NLP solver may output trajectories that violate the dynamic and safety constraints, causing catastrophic damage to safety-critical systems.
+It also lacks the ability to verify the optimality of the sampled trajectories.
+While efforts have been made to integrate safety constraints into generative models (Li et al., 2023; Chang et al., 2023; Yang et al., 2023; Xiao et al., 2023; Power et al., 2023; Botteghi et al., 2023), it is still challenging to achieve solution quality, diversity, feasibility, and computational efficiency at the same time.
+In this paper, we propose DiffuSolve: a diffusion model-based solver for non-convex trajectory optimization with improved efficiency, diversity, and robustness.
+Our contribution is two-fold: First, we train a conditional diffusion model to learn the complex solution distribution of non-convex optimization problems, and efficiently sample a set of diverse and high-quality trajectory predictions.
+These samples are then used as warm starts for an NLP solver to derive locally optimal and feasible solutions, ensuring the robustness of DiffuSolve.
+Second, to mitigate the effect of diffusion samples that violate the problem constraints, we developed DiffuSolve+ with a novel constrained diffusion model that incorporates an additional constraint violation loss in the training process.
+DiffuSolve has the following appealing properties: (a).
 Computational Efficiency.
-Trained with offline solution data, the diffusion model can predict high-quality solutions that are close to local optima and warm-start the trajectory optimization.
-This significantly accelerates the solver convergence with considerable computational time savings.
+The diffusion model samples high-quality solutions that are close to local optima as warm starts, from which the NLP solver can quickly converge.
 (b).
-Robustness (Constraint Satisfaction).
-We used a constrained diffusion model specifically designed to produce samples with minimal constraint violations.
-These sampled trajectories are then further refined by a numerical solver to enhance the feasibility and optimality.
+Robustness and Diversity.
+A constrained diffusion model produces diverse samples with minimal constraint violations, further refined by an NLP solver for enhanced feasibility and optimality.
 (c).
 Automatic Data Generation.
-With a numerical solver, we can automatically generate diverse problems and collect corresponding solutions, which allows for on-demand data augmentation whenever required.
+With an NLP solver, we can generate diverse problems and collect corresponding solutions ourselves, allowing for on-demand data augmentation whenever required.
 (d).
 Generalization.
-Our framework is suitable for trajectory optimization across various applications with flexible environment settings.
-We demonstrate the efficacy of our framework on three non-convex trajectory optimization problems for open-loop optimal control: tabletop manipulation, two-car reach-avoid, and cislunar low thrust transfer for a spacecraft.
-Our framework improves the efficiency of obtaining solutions compared to the baseline while enhancing constraint satisfaction.
+Our framework is suitable for trajectory optimization across various system dynamics and environment settings, with the potential for optimization problems in other areas like finance (Sambharya et al., 2023), etc.
+We demonstrate the efficacy of our method on three non-convex trajectory optimization problems: two-car reach-avoid, quadrotor navigation, and cislunar low-thrust transfer for a spacecraft.
 2 Related Work
 Machine Learning for Optimization.
-Machine learning methods have been widely used to solve a variety of optimization problems with different techniques, including convex Quadratic Programming (QP) [50, 51, 11, 38], mixed-integer programming [9], combinatorial optimization [43, 23] and black-box optimization [30, 45, 28], etc.
-To improve the constraint satisfaction in optimization problems, neural networks have been incorporated with fully differentiable constraint function [16].
-Conditional variational autoencoder (CVAE) and long short-term memory (LSTM) are proposed to warm-start the general nonlinear trajectory optimization problems [32].
-Another branch of work leverages machine learning methods and pre-collected solution data points for solving partial differential equations (PDEs) [7].
-One of the most pupular work is physics-informed neural networks (PINN) [37], which introduces the solution prediction model with a hybrid training loss for labeled solution prediction and violation of the original PDEs.
+Machine learning methods have been widely used to solve a variety of optimization problems with different techniques, including convex Quadratic Programming (QP) (Zeilinger et al., 2011; Zhang et al., 2019; Chen et al., 2022; Sambharya et al., 2023), mixed-integer programming (Cauligi et al., 2021), combinatorial optimization (Sun & Yang, 2023; Huang et al., 2023) and black-box optimization (Kumar & Levine, 2020; Trabucco et al., 2022; Krishnamoorthy et al., 2023), etc.
+For the Combinatorial Optimization problem, graph-based diffusion models are used to directly generate solutions (Sun & Yang, 2023; Huang et al., 2023).
+To improve the constraint satisfaction in optimization problems, neural networks have been incorporated with fully differentiable constraint function (Donti et al., 2021).
+Conditional variational autoencoder (CVAE) and long short-term memory (LSTM) are proposed to warm-start the general nonlinear trajectory optimization problems (Li et al., 2023).
+Another branch of work leverages machine learning methods and pre-collected solution data points for solving partial differential equations (PDEs) (Cai et al., 2021).
+One of the most pupular work is physics-informed neural networks (PINN) (Raissi et al., 2019), which introduces the solution prediction model with a hybrid training loss for labeled solution prediction and violation of the original PDEs.
 Diffusion Models.
-The diffusion models [39, 41] is able to sample the multi-modal distribution with a stochastic differential equation (SDE).
-The models are further developed into denoising diffusion probabilistic models (DDPM) [21], denoising diffusion implicit models (DDIM) [40] for application of image generation.
-Both classifier-guided diffusion [13] and classifier-free guidance [22] are proposed to achieve conditional generation with diffusion models.
-Given the strong expressiveness of the diffusion model, its application has been extended across domains from image generation to reinforcement learning [46, 14, 15], motion generation [44], robotics [12], etc.
-Diffusion-based Trajectory Generation.
-Diffusion models are becoming popular in robotics for generating controls or trajectories [1, 24, 33, 42, 12, 34, 15].
+The diffusion models (Sohl-Dickstein et al., 2015; Song et al., 2020b) is able to sample the multi-modal distribution with a stochastic differential equation (SDE).
+The models are further developed into denoising diffusion probabilistic models (DDPM) (Ho et al., 2020), denoising diffusion implicit models (DDIM) (Song et al., 2020a), etc., for application of image generation.
+Both classifier-guided diffusion (Dhariwal & Nichol, 2021) and classifier-free guidance (Ho & Salimans, 2022) are proposed to achieve conditional generation with diffusion models.
+Diffusion-based Trajectory Generation with Constraints.
+Diffusion models are becoming popular in robotics for generating controls or trajectories (Ajay et al., 2022; Janner et al., 2022; Liang et al., 2023; Sridhar et al., 2023; Chi et al., 2023; Mishra et al., 2023; Ding et al., 2024).
 Many efforts have been made to improve the diffusion model for creating safe trajectories for safety-critical systems.
-Control barrier functions [48, 6] and collision-avoidance kernels [10] are employed to guide diffusion models to sample feasible trajectory solutions.
-To adapt to new constraints in testing, prior work investigated composing diffusion models to improve the generalizability [36, 49].
-However, solely relying on diffusion-sampled trajectories will inevitably face the robustness issue of constraint violations and hardly achieve locally optimal solutions.
-There has been an effort to combine diffusion models with an inference-based trajectory optimization framework [36, 35], but it is not parallelizable and requires extensive computational effort for complex problems.
-Our work guides the diffusion model toward low constraint violation in the output through a novel loss function and combines it with a fully parallelizable numerical optimization method that scales well.
+Control barrier functions (Xiao et al., 2023; Botteghi et al., 2023) and collision-avoidance kernels (Chang et al., 2023) are employed to guide diffusion models to sample feasible trajectory solutions.
+To adapt to new constraints in testing, prior work investigated composing diffusion models to improve the generalizability (Power et al., 2023; Yang et al., 2023).
+Different from existing work, our paper focuses on high-dimensional, highly nonlinear trajectory optimization problems with many local optima, and aims to find diverse solutions efficiently.
+Instead of incorporating constraints in the sampling process, our constrained diffusion model in DiffuSolve+ includes a novel loss in training, which can be combined with the current work to further reduce the violation.
+Moreover, DiffuSolve and DiffuSolve+ are equipped with an NLP solver that finetunes the diffusion samples until feasibility and optimality are obtained at its best.
 3 Preliminaries
 Non-Convex Trajectory Optimization.
-We consider a general open-loop optimal control problem: assume a known agent’s dynamical model, the predefined initial configuration, and the obstacle settings, we aim to plan an optimal trajectory and controls for the agent to reach a target while satisfying the dynamics and safety constraints.
-With the direct transcription [4], this trajectory optimization problem can be discretized as the following nonlinear program $\mathcal{P}_{y}$:
+We consider a general open-loop trajectory optimization problem: assume a known agent’s dynamical model, the predefined initial configuration, and the obstacle settings, we aim to plan an optimal trajectory and controls for the agent to reach a target while satisfying the dynamics and safety constraints.
+With the direct transcription (Betts, 1998), this trajectory optimization problem can be discretized as the following nonlinear program $\mathcal{P}_{y}$:
 […]
@@
 (1) typically presents a non-convex optimization problem with multiple local optima.
-To solve this problem $\mathcal{P}_{y}$, a gradient-based numerical solver $\pi$ can be used.
-A local optimum $x^{*}$ to $\mathcal{P}_{y}$ is verified by the solver $\pi$ that satisfies the KKT conditions (i.e., no constraint violation).
+To solve this problem $\mathcal{P}_{y}$, a gradient-based NLP solver $\pi$ can be used.
+A local optimum (feasible) $x^{*}$ to $\mathcal{P}_{y}$ is verified by the solver $\pi$ that satisfies the KKT conditions with no constrained violation.
 To find different $x^{*}$, the user first provides multiple initial guesses $x^{0}$ for the decision variables $x$ to the solver $\pi$, typically drawn from a uniform distribution or some heuristic approach.
 Then the solver $\pi$ performs an iterative, gradient-based optimization until a local optimum $x^{*}$ is found.
-The popular methods to conduct such gradient-based optimization in non-convex problems include the Interior Point Method (IPM) [47] and the Sequential Quadratic Programming (SQP) [5].
+The popular methods to conduct such gradient-based optimization in non-convex problems include the Interior Point Method (IPM) (Wright, 1997) and the Sequential Quadratic Programming (SQP) (Boggs & Tolle, 1995).
 Diffusion Probabilistic Model.
-As expressive generative models, the diffusion probabilistic models [39, 21, 41] generate samples from random noises through an iterative denoising process.
+As expressive generative models, the diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Song et al., 2020b) generate samples from random noises through an iterative denoising process.
 The forward process (i.e., diffusion process) iteratively adds Gaussian noise to sample $x_{k}$ at the $k$-th step according to the following equation:
@@
 where $\widebar{\alpha}_{k}=\prod_{k^{\prime}=1}^{K}(1-\beta_{k^{\prime}})$ and $\varepsilon\sim\mathcal{N}(0,\mathbf{I})$ is a random sample from a standard Gaussian.
-The data generation process is the reverse denoising process that transforms random noise to data sample:
+The data generation process is the reverse denoising process that transforms random noise into data sample:
 $p_{\theta}(x_{k-1}|x_{k})=\mathcal{N}\big{(}x_{k-1};\mu_{\theta}(x_{k}),\sigma_{k}\mathbf{I}\big{)},\quad 1\leq k\leq K$ (4)
@@
 For conditional generation modeling, a conditional variable $y$ is added to both processes $q(x_{k+1}|x_{k},y)$ and $p_{\theta}(x_{k-1}|x_{k},y)$, respectively.
-The classifier-free guidance [22] can be applied to further promote the conditional information, which learns both conditioned and unconditioned noise predictors as $\varepsilon_{\theta}(x_{k},k,y)$ and $\varepsilon_{\theta}(x_{k},k,\varnothing)$.
+The classifier-free guidance (Ho & Salimans, 2022) can be applied to further promote the conditional information, which learns both conditioned and unconditioned noise predictors as $\varepsilon_{\theta}(x_{k},k,y)$ and $\varepsilon_{\theta}(x_{k},k,\varnothing)$.
 Specifically, $\varepsilon_{\theta}$ is optimized with the following loss function:
@@
 4 Methodology
-In this section, we first proposed the general framework for efficient and robust non-convex trajectory optimization using diffusion models and numerical solvers.
-Then we introduce a novel constrained diffusion model specifically designed to guide the diffusion models to generate feasible solutions.
-4.1 A General Trajectory Optimization Framework
-Figure 2: The proposed general framework for non-convex trajectory optimization with a constrained diffusion model and a numerical solver.
+We first propose DiffuSolve, a general framework for efficient and robust non-convex trajectory optimization using diffusion models and NLP solvers.
+Then we introduce a novel constrained diffusion model specifically designed to guide the diffusion models to generate feasible solutions, which is included in DiffuSolve+.
+4.1 DiffuSolve Framework
+Figure 2: The DiffuSolve framework for solving non-convex trajectory optimization problems.
 It has been observed that for similar trajectory optimization problems $\mathcal{P}_{y}$ in Eq.
-(1), their solutions $x^{*}$ often exhibit similar structures [2, 32].
+(1), their solutions $x^{*}$ often exhibit similar structures (Amos et al., 2023; Li et al., 2023).
 Similar problems are those $\mathcal{P}_{y}$ with the same objective $J$ and constraints $g,h$, but with different parameters $y$.
-Thus, the key insight of our framework is to use a diffusion model to learn $p(x^{*}|y)$, the conditional probability distribution of locally optimal solutions, based on pre-solved similar problems.
+Thus, the key insight of DiffuSolve is to use a diffusion model to learn $p(x^{*}|y)$, the conditional probability distribution of locally optimal solutions based on pre-solved similar problems.
 Then the diffusion model can generalize and predict diverse solutions $\tilde{x}_{\text{new}}^{*}$ to new scenarios where the condition $y_{\text{new}}$ value is unseen.
 However, diffusion models inevitably make prediction errors, which can result in constraint violations.
-To address this, we use diffusion samples $\tilde{x}_{\text{new}}^{*}$ as initial guesses for the numerical solver, which fine-tunes them to derive the final solutions $x_{\text{new}}^{*}$, increasing both feasibility and optimality.
+To address this, we use diffusion samples $\tilde{x}_{\text{new}}^{*}$ as initial guesses for the NLP solver, which fine-tunes them to derive the final solutions $x_{\text{new}}^{*}$, increasing both feasibility and optimality.
 This approach acts like a safety net, enhancing the robustness of the solutions.
-We illustrate our proposed data-driven trajectory optimization framework with diffusion models in Fig.
+We illustrate our proposed DiffuSolve framework in Fig.
 2.
-In the offline data generation process, we use a numerical solver $\pi$ to collect locally optimal solutions $x^{*}$ for various problem instances $\mathcal{P}_{y}$ with uniformly sampled initial guesses.
+In the offline data generation process, we use an NLP solver $\pi$ to collect locally optimal and feasible solutions $x^{*}$ for various problem instances $\mathcal{P}_{y}$ with uniformly sampled initial guesses.
 The solutions with sub-optimal objective values are filtered out manually and the remaining dataset is used as the training data.
-We train a conditional diffusion model to learn $p(x^{*}|y)$ with classifier-free guidance [22].
+For DiffuSolve training, we use a conditional diffusion model to learn $p(x^{*}|y)$ with classifier-free guidance (Ho & Salimans, 2022), with the loss function in Eq.
+5.
 In the online testing process when a new problem $\mathcal{P}_{y_{\text{new}}}$ is presented with a priori unknown value $y_{\text{new}}$, we first use the diffusion model to predict diverse solution candidates $\tilde{x}_{\text{new}}^{*}$ as initial guesses.
-Then the numerical solver $\pi$ only requires minor adjustments to derive the final solutions $x_{\text{new}}^{*}$ with improved feasibility and optimality.
+Then the NLP solver $\pi$ only requires minor adjustments to derive the final solutions $x_{\text{new}}^{*}$ with improved feasibility and optimality.
 This warm-starting approach is fully parallelizable for both diffusion sampling and the solving process of $\pi$.
-4.2 Constrained Diffusion Model
-The original diffusion models (DMs) aim to learn the distribution of locally optimal solutions $p(x^{*}|y)$ through the loss function in Eq.
+4.2 DiffuSolve+ with Constrained Diffusion Model
+The original diffusion models in DiffuSolve aim to learn the distribution of locally optimal solutions $p(x|y),x\in\mathcal{X},y\in\mathcal{Y}$ with $\mathcal{X}$ as the variable space and $\mathcal{Y}$ as the problem parameter space.
+DiffuSolve is optimized through the loss function as Eq.
 (5) without knowing any constraint information.
 As a result, the sampled solutions could be close to the groundtruth but still largely violate the constraints.
-To address this issue, we propose a novel constrained diffusion model (CDM) to minimize the constraint violations in the sampled trajectories.
-Inspired by the equation loss in PINN [37], we define a violation function $V(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for differentiable constraints in Eq.
+To address this issue, we propose an improved DiffuSolve+ method with a constrained diffusion model (CDM) to minimize the constraint violations in the sampled trajectories.
+Inspired by the equation loss in PINN (Raissi et al., 2019), we define a violation function $V(x,y):\mathcal{X}\times\mathcal{Y}\rightarrow\mathbb{R}$ for differentiable constraints in Eq.
 (1):
 $V=\sum_{i=1}^{l}\max(g_{i},0)+\sum_{j=1}^{m}|h_{j}|,$ (6)
-where $V(\cdot,\cdot)\in C^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ (differentiable except at $g_{i}=0$ or $h_{j}=0$) maps from a sample $x$ and parameter $y$ and outputs the total constraint violation as a scalar value.
+where $V(\cdot,\cdot)\in PC^{1}(\mathbb{R}^{n},\mathbb{R}^{k};\mathbb{R})$ (piecewise differentiable except at $g_{i}=0$ or $h_{j}=0$) maps from a sample $x$ and parameter $y$ and outputs the total constraint violation as a scalar value.
 In fact, each constraint term can be further customized with different weights that allow different treatments.
@@
 We introduce the newly proposed loss function and the corresponding training process as follows.
-CDM Training.
-We define the locally optimal solution $x^{*}=x_{0}$ as initial sample in the diffusion process with diffusion step $k=0$.
+DiffuSolve+ Training.
+For the diffusion process, we try to generate samples from solution set $x_{0}\in\mathcal{X}$.
 By reformulating the forward sampling process as in Eq.
 (3), we have $x_{0}=\frac{x_{k}-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon}{\sqrt{\widebar{\alpha}_{k}}},\varepsilon\sim\mathcal{N}(0,\mathbf{I})$.
-The noise $\varepsilon$ is approximated as $\varepsilon_{\theta}(x_{k},k,y)$ with neural-network parameterization.
-Therefore, we can predict initial sample as $\widehat{x}_{0}=\frac{x_{k}(x_{0},\varepsilon)-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon_{\theta}(x_{k},k,y)}{\sqrt{\widebar{\alpha}_{k}}}$.
-Then we introduce a constraint violation loss $\mathcal{L}_{\text{vio}}$ on the predicted $\widehat{x}_{0}$ from $x_{k}$, $\forall k\in[K]$:
+The noise $\varepsilon$ is approximated as the neural-network parameterized model $\varepsilon_{\theta}(x_{k},k,y)$ for condition $y$ at timestep $k$.
+Replacing it into previous formula, the clean sample at timestep $k=0$ is predicted as $\widehat{x}_{0}=\frac{x_{k}(x_{0},\varepsilon)-\sqrt{1-\widebar{\alpha}_{k}}\varepsilon_{\theta}(x_{k},k,y)}{\sqrt{\widebar{\alpha}_{k}}}$.
+The benefits of direct $\widehat{x}_{0}$ prediction is that it does not need iterative diffusion sampling thus avoids gradient backpropagation through multi-step sampling.
+For constrained diffusion generation with DiffuSolve+, we introduce a constraint violation loss $\mathcal{L}_{\text{vio}}$ on the predicted $\widehat{x}_{0}$ from $x_{k}$, $\forall k\in[K]$:
 $\mathcal{L}_{\text{vio}}$ $=\mathbb{E}_{(x_{0},y)\sim\mathcal{X}\times\mathcal{Y},k\in[K],\varepsilon\sim\mathcal{N}(0,\mathbf{I})}\bigg{[}\frac{1}{k}V\Big{(}\text{clip}\big{(}\widehat{x}_{0}(x_{k},\varepsilon_{\theta}),x_{0}^{\text{min}},x_{0}^{\text{max}}\big{)},y\Big{)}\bigg{]}$ (7)
-where $x_{0}^{\text{min}}$ and $x_{0}^{\text{max}}$ are the lower and upper bound for $x_{0}\in\mathcal{X}$.
+where the clip function preserves the range of $\widehat{x}_{0}$ to be within the lower $x_{0}^{\text{min}}$ and upper bound $x_{0}^{\text{max}}$ for $x_{0}\in\mathcal{X}$, ensuring the valid calculation of violation function $V(\cdot,\cdot)$.
 The reason for the clipping is that at the early phase of the training, the parameterized $\varepsilon_{\theta}$ as an approximation of $\varepsilon$ is not accurate thus the predicted $\widehat{x}_{0}$ may not lie on the original range of the data $x_{0}$.
-Therefore, we clip $\widehat{x}_{0}$ to maintain its range, which guarantees the valid calculation of violation function $V(\cdot,\cdot)$.
-The scheduling factor $\frac{1}{k}$ regulates the scale of the violation function $V$ for different $x_{k},k\in[K]$.
-The reason is that the predicted $\widehat{x}_{0}$ will be more noisy when $k$ is large therefore the estimation of violation function can be less reliable.
+The scheduling factor $\frac{1}{k}$ regulates the scale of the violation function $V$ for different $x_{k},k\in[K]$, since the predicted $\widehat{x}_{0}$ will be more noisy when $k$ is large and the estimation of violation function can be less reliable.
 This may not be a theoretically principled choice but is found effective and convenient in our experiments.
@@
 An intuitive interpretation of the additional loss term is that it guides the diffusion model training with the gradients of minimizing the violation value for each denoising step.
-Finally, our proposed constrained diffusion training loss is a combination of the original diffusion model training loss and the above violation loss:
+Finally, our proposed constrained diffusion training loss is a combination of the original diffusion model training loss in DiffuSolve as Eq. 5 and the above violation loss:
 $\mathcal{L}=\mathcal{L}_{\text{diff}}+\lambda\mathcal{L}_{\text{vio}}$ (8)
@@
 $\widehat{\varepsilon}_{\theta}\leftarrow c\cdot\varepsilon_{\theta}(x_{k},k,y)+(1-c)\cdot\varepsilon_{\theta}(x_{k},k,\varnothing)$ (9)
-Then we plug $\widehat{\varepsilon}_{\theta}$ into $\varepsilon_{\theta}$ in the sampling process [21, 22], and through iterative sampling $x_{k-1}$ from $x_{k}(k=K,\dots,1)$ with initial random noise $x_{K}\sim\mathcal{N}(0,\mathbf{I})$, it produces the generated samples $x_{0}$.
-This diffusion samples serve as the predicted initial guess in Sec.
+Then we plug $\widehat{\varepsilon}_{\theta}$ into $\varepsilon_{\theta}$ in the sampling process (Ho et al., 2020; Ho & Salimans, 2022), and through iterative sampling $x_{k-1}$ from $x_{k}(k=K,\dots,1)$ with initial random noise $x_{K}\sim\mathcal{N}(0,\mathbf{I})$, it produces the generated samples $x_{0}$.
+The diffusion sample serves as the predicted initial guess in Sec.
 4.1 for a new task: $\tilde{x}^{*}_{\text{new}}:=x_{0}$.
 5 Experiment
-In this section, we evaluate the proposed framework combining diffusion models and numerical solvers on three trajectory optimization tasks, and compare with several baselines.
-Our methods include two DMs: the vanilla diffusion model (DM) and the constrained diffusion model (CDM) as proposed in Sec.
-4.2.
-The experiments involve three stages: 1.
-Collect solutions provided by numerical solvers on each task domain; 2.
-For data-driven methods, train the models to fit the solution dataset; 3.
-For unseen new tasks within each domain, sample initial solutions from models and feed into numerical solvers for deriving final solutions.
+In this section, we first evaluate DiffuSolve with a vanilla diffusion model (DM) on three high-dimensional and highly nonlinear trajectory optimization tasks.
+Secondly, we present the results of DiffuSolve+ with our novel constrained diffusion model (CDM).
+As a standalone method, the CDM reduces problem constraint violation of the samples compared to DM.
+Within DiffuSolve+, CDM helps further improve the efficiency of obtaining locally optimal and feasible solutions compared to DiffuSolve.
+The experiments mainly involve three stages: 1.
+Use an NLP solver to collect solutions on each task domain; 2.
+Train the data-driven models to fit the solution dataset; 3.
+For unseen new tasks within each domain, sample the initial guesses from each model that warm-start the solver to derive final solutions.
 The details of data collection and model training refer to Appendix A.
-Tasks.
-Our experiments are conducted on three trajectory optimization tasks across different domains with examples, as shown in Fig.
-1.
-With an increasing complexity (non-convexity) of the problem, we introduce: (i).
-tabletop manipulation for object relocation with a robotic arm (Sec.
-5.1); (ii).
-the two-car reach-avoid problem as a two-player motion planning game (Sec.
-5.2); and (iii) the cislunar transfer problem for optimal control of the spacecraft (Sec.
-5.3).
-Each task has various DoF, which is derived with the number of control variables at each time step multiplied with the number of discretized steps along the trajectory.
-The tabletop manipulation has a higher DoF due to a larger number of time stesp ($T=80$), however, the task is less complex as the cislunar transfer, which has more complicated dynamics thus higher non-convexity.
-More details for each task refer to Appendix A.
-Numerical Solvers.
-We assume fully known dynamics and environment for the open-loop optimal control problem.
-To solve the tabletop manipulation and two-car reach-avoid problem, we formulate and solve the problem with Sparse Nonlinear OPTimizer (SNOPT) [19] as the solver $\pi$, which uses the SQP method and supports warm-starting.
-For the cislunar transfer, we formulate and solve the problem with pydylan, a Python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] together with SNOPT.
-The detailed problem formulation and solving for each task are included in Appendix A.
-Baselines.
-For comparison with the proposed framework, we introduce the following baselines:
-  • Uniform: This method solely uses a numerical solver with uniformly sampled initial guesses $x\in\mathcal{X}$ for the non-convex trajectory optimization problem.
-It solves problems from scratch without leveraging any problem-specific prior information.
-  • Optimal Uniform: This method uniformly samples from collected locally optimal solutions $\{x^{*}\}$ as initial guesses for the numerical solver, without using any machine learning method to approximate the solution distribution.
-  • CVAE_LSTM: This method [32] combines the convolutional variational auto-encoder (CVAE) with long short-term memory (LSTM) to learn the conditional distribution of locally optimal solutions.
-Specifically, a CVAE with Gaussian Mixture Model (GMM) prior is first used to sample the non-control variables, e.g.
-time or mass variables, and then an LSTM is adapted to sample the control sequence.
-Although there exist other graph-based path planning methods like Rapidly-exploring Random Trees (RRT) [31], A [20], etc., they can not account for both complex dynamical and environment constraints and flexible optimization objectives in the optimal control problem.
-Consequently, the comparison against them is not provided in our experiments.
-Diversity Measure.
-Apart from the time consumption for deriving feasible solutions, we also apply diversity measure for each method since there exist multiple local optima for each task.
-Specifically, the Vendi Score [18] is applied to measure the diversity of the model outputs.
-The similarity kernel we choose is $k(x,y)=e^{-||x-y||}$.
-5.1 Tabletop Manipulation
-Task Setting.
-For a robotic arm manipulating objects on the tabletop as shown in Fig.
-7 in the Appendix, it is important to plan a desired trajectory with goal-reaching and obstacle avoidance.
-We define the states of the system to be the location of the gripper: $(p_{x},p_{y})\in[0,10]^{2}$.
-In a simplified simulation setting, the movement of the gripper is modeled with linear dynamics $\dot{p}_{x}=v_{x},\dot{p}_{y}=v_{y}$, where $u=(v_{x},v_{y}),v_{x},v_{y}\in[-1,1]$ are two-dimensional speed control variable.
-In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles.
-Figure 3: Raw trajectory prediction output and the solver output for the tabletop manipulation (top row) and two-car reach-avoid problem (bottom row).
-From left to right, uniform sampling, CVAE_LSTM, DM, CDM and solver refined locally optimal solution from CDM sampled initial guess.
-Results.
-In the top row of Fig.
-3, our proposed constrained diffuse model generates a nearly collision-free trajectory to reach the goal, which is very similar to the final refined solution from the solver.
-The DM tries to navigate through the cluttered environment but chooses a path that inevitably collides with obstacles.
-The baseline CVAE_LSTM is not able to account for both goal-reaching and obstacle avoidance.
-This can be further verified by checking the constraint violation in Table 2, where both DM and CDM achieve significant reduction on constraint violation.
-CDM has the lowest violation, which justifies the effectiveness of the additional constraint violation loss for training.
-When we use our proposed models to warm-start the solver $\pi$, only minor refinement is needed to derive locally optimal trajectories.
-We compute the ratio of locally optimal solutions obtained from 600 initial guesses and present the corresponding time consumption (involving both sampling time from the model and numerical solver time) statistics in Table 1.
-As demonstrated in Table 1, the CDM with the solver is 29.6 times faster than the uniform method based on the top $25\%$-quantile of the computational time for deriving the locally optimal solutions.
-This is further demonstrated in Fig.
-LABEL:Fig:tabletop_time_histogram with the solving time histogram of locally optimal solutions, where the CDM with the solver starts to obtain many solutions in a shorter solving time.
-Then Vendi score result in Table 1 verifies the diversity of our model output, where uniform samples intrinsically have the highest diversity.
-Thus both efficiency and solution quality are achieved with our framework.
-5.2 Two-Car Reach-Avoid
-Task settings.
-The two-car reach-avoid problem is a two-player game where each car tries to find a minimum-time path to reach the goal while avoiding collision with another car and obstacles simultaneously.
+5.1 Task setup
+Our experiments are conducted on three complex non-convex trajectory optimization tasks as illustrated in Figure 1.
+These tasks span diverse domains and present unique challenges: (i).
+Two-Car Reach-Avoid: Multi-agent game-theoretical problem is difficult to scale and features multiple equilibria.
+(ii).
+Quadrotor Navigation: High-dimensional nonlinear dynamical systems lead to computational inefficiency.
+(iii).
+Cislunar Transfer: The highly nonlinear chaotic dynamical system makes spacecraft trajectory optimization time-consuming with numerous local optima.
+The detailed dynamical settings for each task are as follows.
+The numerical settings can be seen in Appendix A.
+Two-Car Reach-Avoid.
+This is a two-player game inspired by (Chen et al., 2018).
+Each car tries to find a minimum-time path to reach the goal while avoiding collision with another car and obstacles, shown in Figure 1.
 Specifically, both car1 and car2 follows the same dynamics as follows:
@@
 Each car has two-dimensional controls $u^{1}=(a^{1},\omega^{1}),u^{2}=(a^{2},\omega^{2})$, where $a\in[-1,1]$ is the acceleration and $\omega\in[-1,1]$ is the angular speed.
-The task setup can be seen in Fig.
-1.
-Table 1: The ratio of locally optimal solutions, the statistics (mean, standard deviation, quantile and median) of total time consumption including both model sampling time and solver time, and Vendi scores for diversity of sampled trajectories before feeding to the solver.
-Task Method Ratio Mean($\pm$±STD) 25%-Quantile Median Vendi Score
-Tabletop CDM (Ours) 59% 7.29 $\pm$ 16.57 0.50 1.26 5190.70
-DM (Ours) 60% 7.51 $\pm$ 17.50 0.73 1.37 5306.18
-CVAE_LSTM 53% 8.53 $\pm$ 20.87 1.10 1.55 516.61
-Uniform 62% 31.65 $\pm$ 24.55 14.81 22.72 5917.73
-Optimal Uniform 77% 12.80 $\pm$ 17.70 4.11 7.17 3708.58
-Two-car CDM (Ours) 95% 17.86 $\pm$ 14.28 7.68 13.83 3873.60
-DM (Ours) 95% 18.82 $\pm$ 14.33 8.77 15.61 3574.66
-CVAE_LSTM 63% 36.24 $\pm$ 20.55 19.45 33.26 1162.72
-Uniform 64% 46.17 $\pm$ 20.63 29.62 43.54 5256.75
-Optimal Uniform 83% 25.15 $\pm$ 19.75 10.82 20.89 3708.58
-Cislunar DM (Ours) 54% 50.81 $\pm$ 70.30 9.97 26.68 5999.96
-CVAE_LSTM 30% 141.31 $\pm$ 125.60 38.41 106.48 1403.38
-Uniform 18% 199.34 $\pm$ 116.59 114.02 174.35 6000
-Optimal Uniform 23% 177.35 $\pm$ 123.22 70.03 153.97 5918.08
-Table 2: The constraint violation value for generated samples before feeding to the solver.
-Task Method Mean($\pm$±STD)
-Tabletop CDM (Ours) 4.60 $\pm$ 5.38
-DM (Ours) 4.73 $\pm$ 5.73
-CVAE_LSTM 9.11 $\pm$ 8.16
-Uniform 32.20 $\pm$ 3.73
-Optimal Uniform 65.44 $\pm$ 40.22
-Two-car CDM (Ours) 3.00 $\pm$ 15.36
-DM (Ours) 10.99 $\pm$ 35.59
-CVAE_LSTM 282.72 $\pm$ 186.40
-Uniform 545.66 $\pm$ 294.71
-Optimal Uniform 3.27 $\pm$ 3.57
-Results.
-In the bottom row of Fig.
-3, the proposed CDM generates the best prediction of the trajectory solution conditioned on obstacle information $y$, which is very close to the final locally optimal solutions refined by the solver.
-The unconstrained DM samples reasonable trajectory solution prediction, but with much higher constraint violation.
-The baseline CVAE_LSTM cannot generate solutions toward goal-reaching.
-Uniformly sampled initial guesses are usually far from a constraint-satisfying trajectory.
-We also evaluate the solving time of obtaining those locally optimal solutions $x^{*}$ from the solver $\pi$ for these five methods in Table 1, and visualize the corresponding histogram in Fig.
-LABEL:Fig:car_solving_time_histogram.
-Our proposed CDMs obtained the greatest number of locally optimal solutions from 600 samples as initializations.
-It is also 3.9 times faster than the uniform method according to the $25\%$-quantile of the computational time for deriving the locally optimal solutions.
-To further compare the CDM and the unconstrained DM, we also measure the constraint violation (as defined by Eq.
-(6)) of all the model outputs in Table 2.
-Our CDM achieves the lowest constraint violation among all data-driven models.
-Note that the optimal uniform also achieves small violation because in this task the goal conditions are fixed, the pre-collected solutions will only incur obstacle collision costs which are distance to the obstacle boundary and are inherently small.
-The Vendi score in Table 1 shows the superiority of using diffusion models to generate diverse outputs compared to the baseline CVAE_LSTM.
-We also visualize the qualitatively different trajectory solution with the proposed CDM and the corresponding solver in Figure 4.
-Figure 4: Visualization of four solutions for one two-car reach-avoid problem, with CDM predicted initial guesses refined by the solver $\pi$.
-5.3 Cislunar Transfer
-Task settings.
-In this task, we consider a minimum-fuel cislunar transfer with the dynamical model to be a Circular Restricted Three-Body Problem (CR3BP) [27].
-Assuming the mass of the spacecraft is negligible, CR3BP describes the equation of motions of a spacecraft under the gravitational force from the sun and moon.
+Quadrotor Navigation.
+This a navigation problem of a 10D quadrotor, inspired by (Herbert et al., 2017).
+The quadrotor needs to find the minimum time to reach the goal position while avoiding obstacles shown in Figure 1, where the goal and obstacles are randomly sampled.
+The quadrotor’s dynamics is defined as follows:
+$\dot{x}$ $=v_{x},$ $\dot{v}_{x}$ $=g\tan\theta_{x},$ $\dot{\theta}_{x}$ $=-d_{1}\theta_{x}+\omega_{x},$ $\dot{\omega}_{x}$ $=-d_{0}\theta_{x}+n_{0}a_{x},$ $\dot{z}$ $=v_{z},$
+$\dot{y}$ $=v_{y},$ $\dot{v}_{y}$ $=g\tan\theta_{y},$ $\dot{\theta}_{y}$ $=-d_{1}\theta_{y}+\omega_{y},$ $\dot{\omega}_{y}$ $=-d_{0}\theta_{y}+n_{0}a_{y},$ $\dot{v}_{z}$ $=K_{T}a_{z}-g$ (10)
+where $x,y,z$ are the position of the quadrotor.
+$\theta_{x}$, $\theta_{y}$ are the pitch and roll angle, and $\omega_{x}$, $\omega_{y}$ are the corresponding rates.
+$v_{x},v_{y},v_{z}$ are the speed of the quadrotor.
+$d_{0},d_{1},K_{T},n_{0}$ are system parameters, and $g=9.81$.
+The controls are $u=(a_{x},a_{y},a_{z})$.
+Cislunar Transfer.
+In this task, we consider a minimum-fuel cislunar low-thrust transfer with the dynamical model to be a Circular Restricted Three-Body Problem (CR3BP) (Koon et al., 2000).
+As shown in Figure 1, the example trajectory (orange) starts from the end of a geostationary transfer spiral (blue) and ends at an arc of the invariant manifold (black).
+Assuming the mass of the spacecraft is negligible, the CR3BP describes the equation of motion of a spacecraft under the gravitational force from the Earth and moon.
 Let $m_{1}$ be the mass of the Earth and $m_{2}$ be the mass of the moon, and $\mu=m_{2}/(m_{1}+m_{2})$, the dynamics of the spacecraft in the CR3BP is as follows:
 $\ddot{x}-2\dot{y}=-\widebar{U}_{x},\quad\ddot{y}+2\dot{x}=-\widebar{U}_{y},\quad\ddot{z}=-\widebar{U}_{z},$
-where $\widebar{U}(r_{1}(x,y,z),r_{2}(x,y,z))=-\frac{1}{2}\left((1-\mu)r_{1}^{2}+\mu r_{2}^{2}\right)-\frac{1-\mu}{r_{1}}-\frac{\mu}{r_{2}}$ represents the effective gravitational potential, $r_{1}=\sqrt{(x+\mu)^{2}+y^{2}+z^{2})}$ and $r_{2}=\sqrt{(x-(1-\mu))^{2}+y^{2}+z^{2})}$ are the distance from the spacecraft to the sun and moon, respectively.
-The objective is to minimize fuel expenditure.
-Results.
-The cislunar transfer is a challenging problem to solve that requires intensive computational time.
-In Table 1, our method with the unconstrained DM has a huge improvement in the average solving time and the number of locally optimal solutions obtained, compared to the baseline.
-It is also 7.0 times faster than the uniform method based on the $25\%$-quantile of computational time for deriving the locally optimal solutions.
-This can be further verified by the solving time histogram in Fig.
-LABEL:Fig:_cr3bp_solving_time_histogram.
-This improvement encourages us to set a short cut-off time for the solver and use more time to sample different initial guesses, which further improves the efficiency of obtaining solutions.
-Finally, the Vendi score shown in Table 1 indicates that our unconstrained DM generates much more diverse predictions than CVAE_LSTM.
-Since the cislunar transfer task uses a complex numerical integration method for the dynamics constraint which requires extra effort to incorporate it into the neural network, we will leave the CDM for this task as future work.
+where $\widebar{U}(r_{1}(x,y,z),r_{2}(x,y,z))=-\frac{1}{2}\left((1-\mu)r_{1}^{2}+\mu r_{2}^{2}\right)-\frac{1-\mu}{r_{1}}-\frac{\mu}{r_{2}}$ is the effective gravitational potential, $r_{1}=\sqrt{(x+\mu)^{2}+y^{2}+z^{2})}$ and $r_{2}=\sqrt{(x-(1-\mu))^{2}+y^{2}+z^{2})}$ are the distance from the spacecraft to the sun and moon, respectively.
+This is a chaotic dynamical system.
+5.2 Experiment Setup
+Numerical Solvers.
+We assume fully known dynamics and environment for the open-loop trajectory optimization problem.
+For all three tasks, we formulate and solve the optimization problem with Sparse Nonlinear OPTimizer (SNOPT) (Gill et al., 2005) as the solver $\pi$, which uses the SQP method and supports warm-starting.
+For the cislunar transfer, we adopt pydylan, a Python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) (Beeson et al., 2022) to formulate the CR3BP dynamics in addition to SNOPT.
+The detailed problem formulation and solving for each task are included in Appendix A.
+Baselines.
+For comparison with the proposed DiffuSolve, we introduce the following baselines:
+  • Uniform: This method solely uses an NLP solver with uniformly sampled initial guesses $x\in\mathcal{X}$ for the non-convex trajectory optimization problem.
+It solves problems from scratch without leveraging any problem-specific prior information.
+  • Optimal Uniform: This method uniformly samples from collected locally optimal solutions $\{x^{*}\}$ as initial guesses for the NLP solver.
+It establishes a performance lower bound for data-driven methods since without learning it cannot generalize to new problems.
+  • CVAE_LSTM: This method (Li et al., 2023) combines the convolutional variational auto-encoder (CVAE) with long short-term memory (LSTM) to learn the conditional distribution of locally optimal solutions.
+Specifically, a CVAE with Gaussian Mixture Model (GMM) prior is first used to sample the non-control variables, e.g.
+time or mass variables, and then an LSTM is adapted to sample the control sequence.
+Although there exist graph-based methods for path planning, such as the Rapidly-exploring Random Trees (RRT) (LaValle, 1998), A (Hart et al., 1968), etc., they do not aim to produce optimal solutions, nor handle the case of nonlinear dynamical systems in complex environmentally constrained problems well.
+Thus, the comparison against them is not provided in our experiments.
+Diversity Measure.
+We adopt the Vendi Score (Friedman & Dieng, 2022) as a quantitative metric for measuring the diversity of the model samples. The similarity kernel we choose is $k(x,y)=e^{-||x-y||}$.
+5.3 DiffuSolve Results
+In this section, we evaluate DiffuSolve’s performance with a vanilla diffusion model (DM) against baseline methods, focusing on solution optimality, feasibility, diversity, and computational efficiency.
+5.3.1 Optimality and Feasibility
+We generate 600 samples from each method and use them as warm starts for the NLP solver.
+In Table 1, we present the percentage (ratio) of locally optimal and feasible solutions over 600 samples which are fine-tuned by the NLP solver.
+Note that before finetuning, this ratio is zero for all methods’ raw samples.
+The optimality and feasibility are verified by checking the KKT condition and the constraint violation based on the task setup in Sec.
+5.1.
+As shown in Table 1, no method is able to directly generate locally optimal and feasible solutions.
+When using the samples as warm starts for the NLP solver, our proposed DiffuSolve obtains a greater number of locally optimal and feasible solutions than any non-diffusion methods.
+We also visualize the sampled trajectories and the corresponding solver-derived trajectories in Figure 3.
+For the Uniform and CVAE_LSTM method, the generated trajectories are far from feasible, thus many of them cannot serve as good initial guesses for the NLP solver.
+DiffuSolve already generates near-feasible trajectories and thus makes NLP solver easier to derive locally optimal and feasible solutions.
+Table 1: Percentage (%) (ratio) of locally optimal and feasible solutions over 600 samples.
+Task Method
+DiffuSolve+ DiffuSolve CVAE_LSTM Uniform Optimal Uniform
+Two-car 94.83 94.50 63.17 64.17 83.00
+Quadrotor 100.00 99.83 99.67 99.17 98.50
+Cislunar 54.00 29.17 17.50 23.33
+Note: All values are for ’+ Solver’ (measured on solver outputs with initial guesses from model samples).
+’- Solver’ values (measured on raw model samples) were 0% for all cases and are omitted for brevity.
+Figure 3: In two-car reach-avoid problem: First row: raw samples (no solver); Second row: corresponding solver derived locally optimal and feasible solutions.
+5.3.2 Computational Efficiency
+For the locally optimal and feasible solutions obtained in Table 1, Table 2 presents the computational time statistics, involving both sampling time from the model and the NLP solver time.
+In Table 2, the DiffuSolve with a vanilla DM is 2 $\times$ to 11 $\times$ faster than the uniform method based on the top $25\%$-quantile of the computational time for deriving the locally optimal and feasible solutions.
+Also in the histogram plots in Figure 5, DiffuSolve obtains more solutions than baseline methods within a short period of time.
+These results coincide with Figure 3 that since the DiffuSolve generates samples that are close to local optima, it takes a shorter time for the NLP solver to converge.
+Thus, it motivates us to set a proper cut-off time for the solver in the future, which can allow more initial guesses to be fed.
+This could further improve the efficiency of DiffuSolve to derive locally optimal and feasible solutions.
+Table 2: Computational time (sampling + solving) statistics and Vendi Score.
+Task Method Solving Time Vendi Score
+Mean($\pm$±STD) 25%-Quantile Median
+Two-car DiffuSolve+ (Ours) 17.86 $\pm$ 14.28 7.68 13.83 3873.60
+DiffuSolve (Ours) 18.82 $\pm$ 14.33 8.77 15.61 3574.66
+CVAE_LSTM 36.24 $\pm$ 20.55 19.45 33.26 1162.72
+Uniform 46.17 $\pm$ 20.63 29.62 43.54 5256.75
+Optimal Uniform 25.15 $\pm$ 19.75 10.82 20.89 3708.58
+Quadrotor DiffuSolve+ (Ours) 6.65 $\pm$ 3.55 4.47 6.05 5458.23
+DiffuSolve (Ours) 7.30 $\pm$ 4.79 4.78 6.62 5428.63
+CVAE_LSTM 9.05 $\pm$ 8.30 5.00 6.74 309.18
+Uniform 16.87 $\pm$ 13.60 9.28 12.48 5985.10
+Optimal Uniform 13.94 $\pm$ 12.08 7.66 10.13 5575.83
+Cislunar DiffuSolve (Ours) 50.81 $\pm$ 70.30 9.97 26.68 5999.96
+CVAE_LSTM 141.31 $\pm$ 125.60 38.41 106.48 1403.38
+Uniform 199.34 $\pm$ 116.59 114.02 174.35 6000.00
+Optimal Uniform 177.35 $\pm$ 123.22 70.03 153.97 5918.08
+5.3.3 Diversity
+In this paper, our proposed DiffuSolve aims to learn the solution distribution of the non-convex problem, instead of predicting a single solution.
+For example, for the two-car reach-avoid problem, there are multiple equilibrium points that are locally optimal and feasible.
+In Figure 4, we present 4 qualitatively different solutions with DiffuSolve for each task, showing its ability to generate diverse solutions even within a shorter period of time.
+The Vendi Score is also evaluated on the samples from each method in Table 2.
+While the Uniform method certainly generates the most diverse samples, the DiffuSolve is comparable to Optimal Uniform and outperforms CVAE_LSTM in sample diversity.
+Especially for the quadrotor navigation, though the solving time for CVAE_LSTM is competitive in Table 2, the Vendi Score shows that its generated solutions are not diverse at all.
+Figure 4: Diverse trajectory solutions for two-car reach-avoid, quadrotor navigation and cislunar transfer problem with DiffuSolve method, given the same conditional input.
+5.4 DiffuSolve+ Results
+In this section, we present the results of DiffuSolve+ on the quadrotor navigation and two-car reach-avoid task.
+The proposed CDM is evaluated both as a standalone method for reducing constraint violations and within the DiffuSolve+ framework for improving computational efficiency.
+Since the current cislunar transfer task originally adopted from (Beeson et al., 2022) uses an adaptive-stepsize integrator for the dynamics constraint, it requires extra effort to incorporate it into the neural network and is beyond our scope.
+We leave the CDM for this task as future work.
+5.4.1 Constraint violation
+As a standalone method without the NLP solver, our proposed CDM from Sec.
+4.2 has fewer constraint violations compared to the vanilla DM and CVAE_LSTM, displayed in Table 3.
+For the quadrotor navigation, the CDM is able to directly generate feasible solutions more than any other methods.
+Also in Figure 3 for the two-car reach-avoid problem, the CDM generates a trajectory that is the closest to the locally optimal and feasible solutions - only car 1 cannot reach the goal but it is already near.
+It will be easy for the NLP solver to close this gap within DiffuSolve+.
+Table 3: Left: The constraint violation of the model samples.
+Right: The ratio of feasible solutions.
+Both are measured directly on the model samples without the NLP solver.
+Here DiffuSolve+: constrained diffusion model, DiffuSolve: vanilla diffusion model
+Task Method Constraint Violation (No Solver) Feasible Ratio ‰
+Mean($\pm$±STD) 25%-Quantile Median (No solver)
+Two-car DiffuSolve+ 3.00 $\pm$ 15.36 0.73 1.44 0
+DiffuSolve 10.99 $\pm$ 35.59 1.98 4.52 0
+CVAE_LSTM 282.72 $\pm$ 186.40 147.09 292.90 0
+Quadrotor DiffuSolve+ 3.06 $\pm$ 7.47 0.10 0.38 81.8
+DiffuSolve 3.85 $\pm$ 8.52 0.16 0.63 49.6
+CVAE_LSTM 20.19 $\pm$ 22.32 3.36 12.96 16.3
+5.4.2 Computational Efficiency
+For DiffuSolve+ with the proposed CDM, the computational efficiency is further improved as shown in Table 2.
+The intuition is clear from Figure 3: when the DiffuSolve+ sample is closer to the local optima, it takes the NLP solver an even shorter time to derive the final solution.
+We also present the histogram of the computational time in Figure 5.
+Here we visualize the density of locally optimal and feasible solutions obtained within a time range, where DiffuSolve+ in red is able to obtain more solutions than DiffuSolve within a short period of time from 600 initial guesses.
+(a) Two-car reach-avoid.
+(b) Quadrotor Navigation.
+(c) Cislunar transfer.
+Figure 5: The histogram of computational time (including model sampling time and solver time) for different methods to find locally optimal and feasible solutions.
 6 Conclusion and Discussion
-This paper presents a general and parallelizable framework for non-convex trajectory optimization, using a novel constrained diffusion model combined with numerical solvers.
-It efficiently provides divesrse solutions for problems with mutiple local optima, with less constraint violation, and up to 30 times acceleration.
-It is a general acceleration framework for solving scientific optimization problems.
-Limitations also exist for the current framework.
-Although most constraint functions in trajectory optimization are differentiable (almost everywhere), current CDM does not work for non-differentiable violation functions.
+This paper presents DiffuSolve: a general diffusion model-based solver for non-convex trajectory optimization.
+It can generate locally optimal, feasible, and diverse solutions with high computational efficiency.
+We also propose DiffuSolve+ which includes a novel constrained diffusion model.
+It has fewer constraint violations as a standalone method and further improves the computational efficiency when integrated with DiffuSolve+.
+It is also a general acceleration framework that has the potential to solve optimization problems in other areas beyond trajectory optimization such as finance, computational chemistry, etc.
+Limitations also exist for the current paper.
+We don’t spend extra effort to optimize the diffusion model’s sampling speed.
+As a result, the current DiffusSolve and CDM may not be able to achieve real-time optimization for high-dimensional systems or highly dynamic environments.
+This can be improved with existing acceleration techniques such as stride sampling, DDIM (Song et al., 2020a), etc.
+In addition, although most constraint functions in trajectory optimization are differentiable (almost everywhere), the current CDM does not work for non-differentiable violation functions.
 For non-differentiable constraints, transforming them into differentiable functions or applying approximation with appropriate bump functions or neural networks is possible.
-Another problem is that the diffusion sampling may not be achieved in real time in a highly dynamic environment, which can be improved with stride sampling, DDIM [40], etc.
 Appendix A Implementation and Task Details
 A.1 Model Details
-For both unconstrained and constrained diffusion models, we apply a UNet structure with three hidden layers of 512, 512 and 1024 neurons.
+For both vanilla and constrained diffusion models (DMs), we apply a UNet structure with three hidden layers of 512, 512, and 1024 neurons.
 We also use a fully connected layer of 256 and 512 neurons to embed the conditional input $y$.
 The sampling step is set to be 500.
-Both DMs are trained with 200 epochs using the Adam optimizer [26].
-The constrained and unconstrained DMs usually take 14 hours and 6 hours respectively, as well as 20 GB VRAM to train on one NVIDIA A100 GPU, respectively.
-A.2 Tabletop Manipulation
-Problem formulation and solving.
-In this task $\mathcal{P}_{y}$, we aim to minimize the time for the gripper to reach the goal while avoiding the obstacles.
-The start position is fixed at $(5,5)$.
-The conditional parameter $y$ includes the goal position $p_{\text{goal}}$ randomly chosen from $\{(1,1),(1,9),(9,1),(9,9)\}$, 4 obstacle positions $p_{\text{obs}}$ randomly sampled between the start and goal, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.0]$.
-We discretize this problem into $T$ timesteps where $T=80$.
-The variable to optimize is $x=(t_{\text{final}},u_{1},u_{2},..,u_{T})\in\mathbb{R}^{161}$, and the conditional parameter is $y=(p_{\text{goal}},p_{\text{obs}},r_{\text{obs}})\in\mathbb{R}^{14}$.
-The objective is defined as $J(x;y)=t_{\text{final}}$.
-The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$ and the collision avoidance for obstacles at all time.
-We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
-Data collection and model training.
-In this task, to collect the training data we sample 2700 different $y$, each with 100 uniformly sampled initial guesses.
-Then we use the SNOPT [19] solver for each problem instance to collect locally optimal solutions.
-Finally, we filter out those solutions with objective values greater than 6 s and use the remaining 237k data as training data.
-Within this 237k data, 10% of the data are used as the validation set.
-The whole data collection process takes around 15 hours with 200 AMD EPYC 9654 CPU cores.
-We train each of our proposed constrained and unconstrained DMs and baseline models with 3 different random seeds.
-The models are tested for the warm-starting performance with 600 initial guesses from 3 random seeds across 20 unseen locations of the goal and obstacles, i.e.
-unseen $y$ values.
-[…]
-[…]
-Figure 7: An example of tabletop manipulation task for goal-reaching and obstacle avoidance movement, using a CDM with a numerical solver for planning: (left) starting position of the gripper and the planned trajectory; (right) target position reached by the gripper.
-A.3 Two Car Reach-Avoid
+Both DMs are trained with 200 epochs using the Adam optimizer (Kingma & Ba, 2014).
+The constrained and unconstrained DMs usually take 14 hours and 9 hours respectively as well as 20 GB VRAM to train on one NVIDIA A100 GPU, for every 100k training data with batch size of 512.
+A.2 Two Car Reach-Avoid
 Problem formulation and solving.
@@
 The conditional parameters $y\in\mathbb{R}^{6}$ include the two obstacle positions $p_{\text{obs}}$ randomly sampled from $[2.0,8.0]$, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[0.5,1.5]$.
-We formulate this trajectory optimization problem with a forward shooting method, with the discretized time step $T=40$.
-The variable to optimize is $x=(t_{\text{final}},u^{1}_{1},u^{1}_{2},...,u^{1}_{T},u^{2}_{1},u^{2}_{2},...,u^{2}_{T})$, where $u^{1}_{i}$ and $u^{2}_{i}$ are controls for two cars, respectively.
+We formulate this trajectory optimization problem with a forward shooting method, with the discretized time step $T=20$.
+The variable to optimize for each car is $x=(t_{\text{final}},u^{1}_{1},u^{1}_{2},...,u^{1}_{T},u^{2}_{1},u^{2}_{2},...,u^{2}_{T})\in\mathbb{R}^{41}$, where $u^{1}_{i}$ and $u^{2}_{i}$ are two control variables for each cars.
 The problem parameter is $y=(p_{\text{obs}},r_{\text{obs}})$.
@@
 Data collection and model training.
-In this task, we collect locally optimal solutions $x^{*}$ from 3k different $y$, each with 100 uniformly sampled initial guesses using SNOPT [19].
+In this task, we collect locally optimal solutions $x^{*}$ from 3k different $y$, each with 100 uniformly sampled initial guesses using SNOPT (Gill et al., 2005).
 We filter out those solutions whose objective is higher than 12 s and use the remaining 114k solutions as the training data.
@@
 We train each of our proposed constrained and unconstrained DMs and baseline models with 3 different random seeds and test the warm-starting performance with 600 initial guesses sampled from 3 random seeds across 20 unseen obstacle settings, i.e.
+unseen $y$ values.
+A.3 Quadrotor Navigation
+Problem formulation and solving.
+In this task $\mathcal{P}_{y}$, we aim to minimize the time for the quadrotor to reach the goal while avoiding the obstacles.
+The start position is fixed at $(-12,0,0)$, and the reference goal position is $(12,0,0)$.
+The conditional parameter $y$ includes the perturbation $d_{\text{goal}}$ randomly chosen from $[-2,2]$ that is added to the goal position, 4 obstacle positions $p_{\text{obs}}$ randomly sampled between the start and goal, and the obstacle radius $r_{\text{obs}}$ randomly sampled from $[1.5,3.5]$.
+We discretize this problem into $T$ timesteps where $T=80$.
+The variable to optimize is $x=(t_{\text{final}},u_{1}^{1},..,u_{T}^{1},u_{1}^{2},...u_{T}^{2},u_{1}^{3},...u_{T}^{3})\in\mathbb{R}^{241}$, and the conditional parameter is $y=(d_{\text{goal}},p_{\text{obs}},r_{\text{obs}})\in\mathbb{R}^{17}$.
+The objective is defined as $J(x;y)=t_{\text{final}}$.
+The constraint $g$ includes the goal-reaching constraints at $t_{\text{final}}$ and the collision avoidance for obstacles at all time.
+We use the 4-th order Runge-Kutta (RK4) method to integrate the dynamics and formulate the constraints.
+Data collection and model training.
+In this task, to collect the training data we sample 4000 different $y$, each with 50 uniformly sampled initial guesses.
+Then we use the SNOPT (Gill et al., 2005) solver for each problem instance to collect locally optimal solutions.
+Finally, we filter out those solutions with objective values greater than 4.57 s and use the remaining 179k data as training data.
+Within this 179k data, 10% of the data are used as the validation set.
+The whole data collection process takes around 20 hours with 200 AMD EPYC 9654 CPU cores.
+We train each of our proposed constrained and unconstrained DMs and baseline models with 3 different random seeds.
+The models are tested for the warm-starting performance with 600 initial guesses from 3 random seeds across 20 unseen locations of the goal and obstacles, i.e.
 unseen $y$ values.
@@
 The variable $x$ to optimize is as follows:
-$\bm{x}=(t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}},m_{f},u_{1},u_{2},...,u_{N})$ (10)
+$x=(t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}},m_{f},u_{1},u_{2},...,u_{N})\in\mathbb{R}^{64}$ (11)
 where $t_{\text{burn}},t_{\text{coast}}^{\text{initial}},t_{\text{coast}}^{\text{final}}$ are the effective burning time for the engine, initial and final coast time, $m_{f}$ is the final mass, and $u_{i}$ is a three-dimensional thrust control of the spacecraft.
@@
 The problem parameter $y$ is the maximum allowable thrust for the spacecraft and is bounded by $[0.1,1]$ Newton.
-To integrate the dynamics, we use the Runge-Kutta-Fehlberg 5-4th order (RK54) method [17].
+To integrate the dynamics, we use the Runge-Kutta-Fehlberg 5-4th order (RK54) method (Fehlberg, 1969).
 Data collection and model training.
 To collect the training data, we sample 12 different $y$ values from a grid in $[0.1,1]$ Newton, and for each $y$ we collect 25k locally optimal solutions with $m_{f}\geq 415kg$ total amount of 300k data.
-The problem is solved using pydylan, a python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) [3] and the solver SNOPT [19].
+The problem is solved using pydylan, a python interface of the Dynamically Leveraged Automated (N) Multibody Trajectory Optimization (DyLAN) (Beeson et al., 2022) and the solver SNOPT (Gill et al., 2005).
 Within this 300k data, 10% of the data are used as the validation set.